### v4

# Temporal-Smoothed Incremental Heuristic Dynamic Programming for Cascaded Online Learning Flight Control System

Abstract
Model-free reinforcement learning has demonstrated the capability to adjust control laws using data-constructed policy gradients. Its application to online flight control commonly adopts a cascaded control architecture, which enables efficient policy learning for angle-of-attack tracking using control-surface deflections. However, the tracking performance and system stability are degraded by oscillatory actions. To address this issue, we incorporate policy regularization for temporal action smoothness and a low-pass filter to attenuate both the amplitude and oscillatory frequency of actions. The Fast Fourier Transform (FFT) algorithm is used to analyze action smoothness in the frequency-domain representation. Flight control simulations demonstrate the effectiveness of the two proposed techniques.
Index Terms:
reinforcement learning, symmetry, data augmentation, approximate value iteration, flight control.
I Introduction
Model-free reinforcement learning (RL) has been applied to flight control law optimization without using an accurate system model for state prediction. This approach significantly reduce the effort required to model the complex dynamics of aerial vehicles in control design, as the policy gradient is constructed using measured data. The ADP-based implementation of model-free RL does not require learning a full system model, since only the first-order derivatives of the system dynamics are needed to construct the policy gradients. This motivates the use of a simpler model with fewer parameters than an artificial neural network, which is commonly used in various ADP methods, while still capturing the required first-order derivatives, such as an incremental model, thereby simplifying training. The incremental model is a local linear approximation of a nonlinear system derived using Taylor expansion. The combination with Recursive Least Squares (RLS) identification in ADP framework has led to the development of incremental model-based ADP (IADP) methods, enabling computationally efficient online policy learning. Compared with Action Dependent Heuristic Dynamic Programming (ADHDP), which implicitly approximates the control effectiveness within a state-action ($Q$) function, the RLS algorithm quickly and accurately identifies the local control effectiveness represented in an incremental model, leading to a more reliable construction of the policy gradient, as demonstrated in simple cases [Ref]. For example, the Incremental model-based Heuristic Dynamic Programming (IHDP) requires only the signs of control effectiveness matrix $G$ to represent the gradient signs, representing less model information than other variants of IADP, including incremental model-based dual heuristic programming (IDHP), and incremental model-based global dual heuristic programming (IGDHP).
The theoretical analysis of online ADHDP has been studied in [Ref], where the gradient descent algorithm is used to update critic and actor over time. The weights are proved to be ultimately uniformly boundedness (UUB) to the sub-optimal weights. learning process involves three time steps. The fact is that IHDP updates the critic in the same way as ADHDP, but updates the actor using an estimated value function reconstructed according to the Bellman equation, instead of a state-action ($Q$) function in ADHDP. As a result, the update laws of the actor are different. In this study, we formulate the update laws of IHDP into two time steps and analyze the UUB convergence of the estimated weights considering action smoothness.
IHDP have been applied to online policy optimization for tracking control of pitch rate, angle of attack (AOA), and altitude for the Flying-V, the Cessna Citation business jet, and the F-16 fighter model.
The application of IHDP for AOA tracking control commonly uses a cascaded control structure to regulate angle-of-attack tracking in the outer-loop system and pitch-rate tracking in the inner-loop system, allowing a hierarchical control strategy that is able to explicitly incorporates pitch-rate constraints. As a comparison, the monolithic structure can also constrain the pitch rate through penalization in the reward design, but its does not learn a pitch rate reference, leading to degraded rate control performance. In addition, the rate control performance is sensitive to the weight of the pitch-rate loss term. Therefore, it is important to develop a well-designed pitch-rate reference for ensuring reliable control performance. An illustrative example is backstepping, where a stabilizing pitch-rate reference policy is derived from a nominal model which ensures the convergence of tracking errors, with accompanying adaptive laws to handle model uncertainties.
However, designing a stabilizing pitch-rate reference is challenging in the model-free IHDP framework, since it relies solely on measured data for pitch-rate reference optimization, which usually leads to a destabilizing reference during learning. A challenge lies in the frequently switching tracking errors near zero caused by time-varying nonlinear dynamics and time-varying $\alpha$ reference signals, which is more serious than in linear time-invariant (LTI) systems. These switching tracking errors, in turn, generate switching pitch-rate references through the outer-loop actor, further perturbing the control surface deflections through the inner-loop actor. This switching phenomenon is also observed in sliding mode control systems.
A second factor contributing to state and action oscillations is actor saturation, in which case the actor uses maximum or minimum actions responding to switching near-zero tracking errors, known as bang–bang control. Meanwhile, vanishing policy gradients in the saturated regions of activation functions cause the actor weights to become nearly fixed and difficult to adjust flexibly. To mitigate this problem, discounted learning rates have been used to slow down actor updates over time, reducing the likelihood of entering the saturation region. The disadvantages lie in that discounted learning rate is selected empirically and is sensitive to its effectiveness. Also, it does not completely prevent actor saturation as learning progresses. The other solution is to terminate actor updates by fixing the weights once the required control performance is achieved, either by setting a termination time or a tracking error threshold, a strategy that has not yet been studied.
Conditioning for Action Policy Smoothness (CAPS) is an action-smoothening technique that regularizes actor smoothness in temporal and spatial scales. The first successful application is on offline-learning quadrotor control [Ref], where the temporal smoothness loss significantly reduces the amplitude of switching motor actions. Other applications are seen for Flying-V, and Cessna Citation. In this study, we only consider regularizing temporal smoothness, given the difficulty of constructing a spatial smoothness loss with limited samples in online learning. The low-pass filter has been used to remove high-frequency noise in command signals in command-filtered backstepping, Incremental Nonlinear Dynamic Inversion (INDI). The advantage of a filter is smoothening high-frequency commands by tuning the natural frequency $\omega_{n}$, even the amplitude of the command is small. This compensates the disadvantage of temporal smoothness in removing low-frequency commands.
The main contributions of this study are summarized as:
  • We propose temporal-smoothed IHDP method to enable smooth policy learning in online learning. The estimated weights of the critic and the actor are analyzed to converge to the sub-optimal weights in UUB sense, considering the extra policy gradients from the temporal smoothness loss.
  • The temporal-smoothed IHDP is applied for cascaded online learning flight control, for angle-of-attack tracking.
  • A low-pass filter is applied to reduce the switching of pitch-rate reference and control surface deflection.
  • The termination and restart conditions are designed based on tracking errors over a fixed past time window, enabling adaptive online learning with flexible control of termination and restart.
The remainder of this paper is structured as follows. Section VI presents the aerial vehicle model and its formulation as a single-input–single-output system. Section LABEL:section_optimization_based_control formulates optimization-based tracking control design. Section LABEL:section_online_learning_implementation presents the cascaded online RL framework, followed by introducing two action-smoothing techniques in Section VI-D. Section VII presents flight control simulations with action-smoothing techniques. Section VIII concludes this paper.
II Incremental Model and Online Identification
II-A Reformulation of Incremental Model
The continuous nonlinear system is given as
$\dot{x}(t)=f(x(t),u(t))$ (1)
where the dynamics $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is a smooth function associated with state vector $x\in\mathbb{R}^{n}$ and input vector $u\in\mathbb{R}^{m}$. $n,m$ are positive integers denoting the dimensions of the state and action spaces.
Taking Taylor expansion for $f(x(t),u(t))$ at $t_{0}$:
$f(x(t),u(t))=$ $f(x(t_{0}),u(t_{0}))+F(t_{0})(x(t)-x(t_{0}))+G(t_{0})(u(t)-u(t_{0}))$ (2)
$+O\left[(x(t)-x(t_{0}))^{2},(u(t)-u(t_{0}))^{2}\right]$
where $F(t_{0})=\partial f(x,u)/\partial x|_{x(t_{0}),u(t_{0})}\in\mathbb{R}^{n\times n}$, $G(t_{0})=\partial f(x,u)/\partial u|_{x(t_{0}),u(t_{0})}\in\mathbb{R}^{n\times m}$ are first-order partial derivatives at $t_{0}$. $O(\cdot)$ is the summed higher-order terms.
The summed higher-order terms $O(\cdot)$ have a relatively small effect on the change from $x_{t}$ to $x_{t+1}$ compared with the first-order terms. Moreover, these higher-order derivatives are difficult to identify from data, which poses challenges for practical application of an incremental model that includes them. For these two reasons, we omit $O(\cdot)$ when defining an effective and practical model for the nonlinear system:
$f(x(t),u(t))=$ $f(x(t_{0}),u(t_{0}))+F(t_{0})(x(t)-x(t_{0}))+G(t_{0})(u(t)-u(t_{0}))$ (3)
Using Equation 104:
$\dot{x}(t)=\dot{x}(t_{0})+F(t_{0})\left[x(t)-x(t_{0})\right]+G(t_{0})\left[u(t)-u(t_{0})\right]$ (4)
The time derivatives are approximated by $\dot{x}(t)\approx\frac{x(t+T)-x(t)}{T}$ and $\dot{x}(t_{0})\approx\frac{x(t_{0}+T)-x(t_{0})}{T}$, where $T$ is a constant time step. Then
$\frac{x(t+T)-x(t)}{T}\approx\frac{x(t_{0}+T)-x(t_{0})}{T}+F(t_{0})\left[x(t)-x(t_{0})\right]+G(t_{0})\left[u(t)-u(t_{0})\right]$ (5)
Define a discretization approach over $T$ by setting $t_{0}=(k-1)T$, $x(t_{0})=x((k-1)T)\triangleq x_{k-1}$. Let $x(t)$ denote the next-step state following $x(t_{0})$, then $x(t)=x(t_{0}+T)\triangleq x_{k}$.
A discretized version of system 108 is
$x_{k+1}-x_{k}\approx x_{k}-x_{k-1}+F_{k-1}(x_{k}-x_{k-1})T+G_{k-1}(u_{k}-u_{k-1})T$ (6)
Use denotions $\Delta x(k+1)=x(k+1)-x(k)$, $\Delta x(k)=x(k)-x(k-1)$, $\Delta u(k)=x(k)-x(k-1)$:
$\Delta x_{k+1}=\Delta x_{k}+F_{k-1}\Delta x_{k}T+G_{k-1}\Delta u_{k}T$ (7)
II-B Recursive Least Squares identification
The linear forgetting Recursive Least Squares (RLS) algorithm is applied to identify matrices $F_{t-1},G_{t-1}$ are identified online. Define the augmented state as $\mathbf{x}_{t}=\left[\Delta q_{t},\Delta\delta_{t}\right]^{T}$ and estimated parameter vector as $\hat{\theta}_{t-1}=[\hat{F}_{t-1},\hat{G}_{t-1}]^{T}$. The one-step state prediction is made as $\Delta\hat{\mathbf{x}}_{t+1}^{T}=\mathbf{x}_{t}^{T}\hat{\theta}_{t-1}$. The error between the true state and the predicted state is denoted by $\varepsilon_{t}$ and defined as $\varepsilon_{t}=\Delta\mathbf{x}_{t+1}^{T}-\Delta\hat{\mathbf{x}}_{t+1}^{T}$. The parameter matrix is updated by
$\hat{\theta}_{t}=\hat{\theta}_{t-1}+P_{t}\mathbf{x}_{t}\varepsilon_{t}$ (8)
$P^{-1}_{t}$ $=\alpha P^{-1}_{t-1}+\mathbf{x}_{t}\mathbf{x}^{T}_{t}$ (9)
$=\sum^{t}_{j=1}\alpha^{t-j}\mathbf{x}_{j}\mathbf{x}^{T}_{j}+\alpha^{t}P^{-1}_{0}$
$=\sum^{t}_{j=1}\alpha^{t-j}\mathbf{x}_{j}\mathbf{x}^{T}_{j}+I\alpha^{t}\lambda_{\text{min}}(0)$
The initial auxiliary matrix is set as $P_{0}=I/\lambda_{\text{min}}(0)$, where $0<\lambda_{\text{min}}(0)\ll 1$ is the minimum eigenvalue of $P^{-1}_{0}$. The initial parameter matrix is $\hat{\theta}_{0}$. The forgetting factor is $0\ll\alpha<1$.
III Temporal-smoothed Incremental Model-based Heuristic Dynamic Programming
III-A Problem formulation
Consider a discrete-time nonlinear dynamical system as
$x(t+1)=f\big(x(t),\,u(t)\big),$ (10)
where $x\in\mathbb{R}^{m}$ is the state vector and $u\in\mathbb{R}^{n}$ is the control vector.
The cost-to-go function is expressed as
$J(x(t))=\sum_{i=t}^{\infty}\alpha^{\,i-t}r\big(x(i),u(i)\big),$ (11)
where $\alpha\in(0,1]$ is the discount factor, $r(x(t),u(t))$ is the reward function. We require $r(t)=r(x(t),u(t))$ to be a bounded semidefinite function of the state and control so that the cost function is well defined. One can obtain from (11) that $0=J(t)-r(t)-\alpha J(t+1)$, where $J(t)=J(x(t))$.
III-B Critic network
The critic network is a one-hidden-layer multilayer perceptron (MLP) that takes the input as $y(t)=\big(x_{1}(t),\ldots,x_{m}(t)\big)^{T}$ and the output as $\hat{J}(t)$ to approximate $J(t)$. The number of nodes in hidden layer is $N_{h_{c}}$. The weight between input node $i$ and hidden node $j$ is $\hat{w}^{(1)}_{c,ji}(t)$. The weight between hidden node $j$ and output node is $\hat{w}^{(2)}_{c,j}(t)$.
The input to hidden node $j$ is:
$\sigma_{c,j}(t)=\sum_{i=1}^{m}\hat{w}^{(1)}_{c,ji}(t)\,x_{i}(t)$ (12)
The output of hidden node $j$ is
$\phi_{c,j}(t)=\tanh(\sigma_{c,j}(t))$ (13)
where $\tanh(\cdot)$ is the nonlinear activation function.
The output of the critic is
$\hat{J}(t)=\sum_{j=1}^{N_{h_{c}}}\hat{w}^{(2)}_{c,j}(t)\,\phi_{c,j}(t),$ (14)
III-C Actor network
The actor network is a one-hidden-layer multilayer perceptron (MLP) that takes the input as the state $x(t)=\big(x_{1}(t),\ldots,x_{m}(t)\big)^{T}$, and the output as the action. The number of nodes in hidden layer is $N_{h_{a}}$. The weight between input node $i$ and hidden node $j$ is $\hat{w}^{(1)}_{a,ji}(t)$. The weight between hidden node $j$ and output node $k$ is $\hat{w}^{(2)}_{a,kj}(t)$.
The input to hidden node $j$ is
$\sigma_{a,j}(t)=\sum_{i=1}^{m}\hat{w}^{(1)}_{a,ji}(t)\,x_{i}(t).$ (15)
The output of hidden node $j$ is:
$\phi_{a,j}(t)=\tanh(\sigma_{a,j}(t))$ (16)
The output of the output node $k$ is
$u_{k}(t)=\sum_{j=1}^{N_{h_{a}}}\hat{w}^{(2)}_{a,kj}(t)\,\phi_{a,j}(t)$ (17)
IV Gradient-descent algorithm
IV-A Update of the critic network
The critic network is trained to minimize the objective function $E_{c}(t)=\frac{1}{2}e_{c}^{2}(t)$, where $e_{c}(t)=\hat{J}(t)-r(t)-\alpha\hat{J}(t+1)$ is the Bellman error. This definition in ADHDP [Ref] follows a backforward style due to the convenience of policy improvement using state-action value function.
The weights between the input layer and the hidden layer are updated by
$\hat{w}^{(1)}_{c}(t+1)=\hat{w}^{(1)}_{c}(t+1)+\Delta\hat{w}^{(1)}_{c}(t)$ (18)
where $\hat{w}^{(1)}_{c}\in\mathbb{R}^{N_{h_{c}}\times m}$.
According to the gradient descent:
$\Delta\hat{w}^{(1)}_{c,ji}(t)=l_{c}\left[-\frac{\partial E_{c}(t)}{\partial\hat{w}^{(1)}_{c,ji}(t)}\right]$ (19)
which follows from the chain rule
$\frac{\partial E_{c}(t)}{\partial\hat{w}^{(1)}_{c,ji}(t)}$ $=\frac{\partial E_{c}(t)}{\partial e_{c}(t)}\frac{\partial e_{c}(t)}{\partial\hat{J}(t)}\frac{\partial\hat{J}(t)}{\partial\phi_{c,j}(t)}\frac{\partial\phi_{c,j}(t)}{\partial\sigma_{c,j}(t)}\frac{\partial\sigma_{c,j}(t)}{\partial\hat{w}^{(1)}_{c,ji}(t)}$ (20)
$=e_{c}(t)\,\hat{w}^{(2)}_{c,j}(t)\phi_{c,j}^{\prime}(t)y_{i}(t)$
The weights between the hidden layer and the output layer are updated by
$\hat{w}^{(2)}_{c}(t+1)=\hat{w}^{(2)}_{c}(t+1)+\Delta\hat{w}^{(2)}_{c}(t)$ (21)
where $\hat{w}^{(2)}_{c}\in\mathbb{R}^{1\times N_{h_{c}}}$.
According to the gradient descent:
$\Delta\hat{w}^{(2)}_{c,j}(t)=l_{c}\left[-\frac{\partial E_{c}(t)}{\partial\hat{w}^{(2)}_{c,j}(t)}\right]$ (22)
which follows from the chain rule
$\frac{\partial E_{c}(t)}{\partial\hat{w}^{(2)}_{c,j}(t)}=\frac{\partial E_{c}(t)}{\partial e_{c}(t)}\frac{\partial e_{c}(t)}{\partial\hat{J}(t)}\frac{\partial\hat{J}(t)}{\partial\hat{w}^{(2)}_{c,j}(t)}=e_{c}(t)\,\phi_{c,j}(t)$ (23)
IV-B Update of the actor network
The actor network is trained according to $\hat{J}(t)$, which is reconstructed from the Bellman equation as $\hat{J}(t)=r(t)+\alpha\hat{J}(t+1)$, rather than a state–action value function $\hat{J}(x(t),u(t))$ as in ADHDP [Ref]. (In this paper, we use an importance-weighted training approach). (Let $U_{c}$ denote the desired ultimate objective function. The quadratic error measure to be minimized is $E_{a}(t)=E^{1}_{a}(t)+E^{2}_{a}(t)$, $E^{1}_{a}(t)=\frac{1}{2}e_{a}^{2}(t),E^{2}_{a}(t)=\frac{1}{2}\lambda e^{2}_{s}(t),e_{a}(t)=r(t)+\alpha\hat{J}(t+1)-U_{c}$, $e_{s}(t)=\frac{1}{2}\|u(t)-u(t+1)\|^{2}$. $e_{s}(t)$ is a temporal smoothness measure. In reinforcement learning, success corresponds to an objective that is zero at each time step. For simplicity of derivation, we assume $U_{c}=0$, i.e. $e_{a}(t)=r(t)+\alpha\hat{J}(t+1)$).
The gradient-descent update rule is $\hat{w}_{a}(t+1)=\hat{w}_{a}(t)+\Delta\hat{w}_{a}(t)$, where $\Delta\hat{w}_{a}(t)=l_{a}\left[-\frac{\partial E_{a}(t)}{\partial\hat{w}_{a}(t)}\right],\ l_{a}>0$.
The weights between the input layer and the hidden layer is updated by
$\hat{w}^{(1)}_{a}(t+1)=\hat{w}^{(1)}_{a}(t+1)+\Delta\hat{w}^{(1)}_{a}(t)$ (24)
where $\hat{w}^{(1)}_{c}\in\mathbb{R}^{N_{h_{c}}\times m}$, and the element-wise increment:
$\Delta\hat{w}^{(1)}_{a,ji}(t)=l_{a}\left[-\frac{\partial E^{1}_{a}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}-\frac{\partial E^{2}_{a}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}\right],$ (25)
Applying the chain rule:
$\frac{\partial E^{1}_{a}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$ $=\frac{\partial E^{1}_{a}(t)}{\partial e_{a}(t)}\left[\frac{\partial e_{a}(t)}{\partial u(t)}\right]^{T}\frac{\partial u(t)}{\partial\phi_{a,j}(t)}\frac{\partial\phi_{a,j}(t)}{\partial\sigma_{a,j}(t)}\frac{\partial\sigma_{a,j}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$ (26)
$=\frac{\partial E^{1}_{a}(t)}{\partial e_{a}(t)}\sum^{n}_{k=1}\frac{\partial e_{a}(t)}{\partial u_{k}(t)}\frac{\partial u_{k}(t)}{\partial\phi_{a,j}(t)}\frac{\partial\phi_{a,j}(t)}{\partial\sigma_{a,j}(t)}\frac{\partial\sigma_{a,j}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$
$=\frac{\partial E^{1}_{a}(t)}{\partial e_{a}(t)}\sum^{n}_{k=1}\left(\frac{\partial r(t)}{\partial u(t)}+\alpha\left[\frac{\partial\hat{J}(t+1)}{\partial x(t+1)}\right]^{T}\frac{\partial x(t+1)}{\partial u(t)}\right)_{k}\frac{\partial u_{k}(t)}{\partial\phi_{a,j}(t)}\frac{\partial\phi_{a,j}(t)}{\partial\sigma_{a,j}(t)}\frac{\partial\sigma_{a,j}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$
$=\frac{\partial E^{1}_{a}(t)}{\partial e_{a}(t)}\sum^{n}_{k=1}\left(2Ru(t)+\alpha\left[\frac{\partial\hat{J}(t+1)}{\partial x(t+1)}\right]^{T}G(t-1)\right)_{k}\frac{\partial u_{k}(t)}{\partial\phi_{a,j}(t)}\frac{\partial\phi_{a,j}(t)}{\partial\sigma_{a,j}(t)}\frac{\partial\sigma_{a,j}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$
$=e_{a}(t)\sum^{n}_{k=1}\left(2Ru(t)+\left[\alpha\hat{w}^{(2)}_{c}(t)\phi^{\prime}(x(t+1))\hat{w}^{(1)}_{c}(t)\right]^{T}G(t-1)\right)_{k}\hat{w}^{(2)}_{a,kj}(t)\phi_{a,j}^{\prime}(x(t))x_{i}(t)$
and
$\frac{\partial E^{2}_{a}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$ $=\frac{\partial E^{2}_{a}(t)}{\partial e_{s}(t)}\left[\frac{\partial e_{s}(t)}{\partial u(t)}\right]^{T}\frac{\partial u(t)}{\partial\phi_{a,j}(t)}\frac{\partial\phi_{a,j}(t)}{\partial\sigma_{a,j}(t)}\frac{\partial\sigma_{a,j}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$ (27)
$=\frac{\partial E^{2}_{a}(t)}{\partial e_{s}(t)}\sum^{n}_{k=1}\frac{\partial e_{s}(t)}{\partial u_{k}(t)}\frac{\partial u_{k}(t)}{\partial\phi_{a,j}(t)}\frac{\partial\phi_{a,j}(t)}{\partial\sigma_{a,j}(t)}\frac{\partial\sigma_{a,j}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$
$=e_{s}(t)\sum^{n}_{k=1}(u_{k}(t)-u_{k}(t+1))\hat{w}^{(2)}_{a,kj}(t)\phi_{a,j}^{\prime}(x(t))x_{i}(t)$
The weight between the hidden layer and the output layer is updated by
$\hat{w}^{(2)}_{a}(t+1)=\hat{w}^{(2)}_{a}(t+1)+\Delta\hat{w}^{(2)}_{a}(t)$ (28)
where $\hat{w}^{(2)}_{a}\in\mathbb{R}^{n\times N_{h_{a}}}$, and the element-wise increment:
$\Delta\hat{w}^{(2)}_{a,kj}(t)=l_{a}\left[-\frac{\partial E^{1}_{a}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}-\frac{\partial E^{2}_{a}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}\right]$ (29)
Applying the chain rule:
$\frac{\partial E^{1}_{a}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}$ $=\frac{\partial E^{1}_{a}(t)}{\partial e_{a}(t)}\frac{\partial e_{a}(t)}{\partial u_{k}(t)}\frac{\partial u_{k}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}$ (30)
$=\frac{\partial E^{1}_{a}(t)}{\partial e_{a}(t)}\left(\frac{\partial r(t)}{\partial u(t)}+\alpha\left[\frac{\partial\hat{J}(t+1)}{\partial x(t+1)}\right]^{T}\frac{\partial x(t+1)}{\partial u(t)}\right)_{k}\frac{\partial u_{k}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}$
$=e_{a}(t)\left(2Ru(t)+\alpha\left[\frac{\partial\hat{J}(t+1)}{\partial x(t+1)}\right]^{T}G(t-1)\right)_{k}\phi_{a,j}(t)$
$=e_{a}(t)\left(2Ru(t)+\left[\alpha\hat{w}^{(2)}_{c}(t)\phi^{\prime}(x(t+1))\hat{w}^{(1)}_{c}(t)\right]^{T}G(t-1)\right)_{k}\phi_{a,j}(t)$
and
$\frac{\partial E^{2}_{a}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}$ $=\frac{\partial E^{2}_{a}(t)}{\partial e_{s}(t)}\frac{\partial e_{s}(t)}{\partial u_{k}(t)}\frac{\partial u_{k}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}$ (31)
$=e_{s}(t)(u_{k}(t)-u_{k}(t+1))\phi_{a,j}(t)$
Remark 1.
The essence of IHDP is abstracting the control effectiveness matrix of the dynamical system G(t1)G(t-1) from a state-action function in Action Dependent Heuristic Dynamic Programming (ADHDP) framework, to improve the learning accuracy and efficiency. The matrix $G$ is identified using Recursive Least Square (RLS) algorithm online.
V Convergence Analysis
V-A Basics of uniformly ultimately bounded stability
Denote the sub-optimal weights by ${w}_{c}^{*},w_{a}^{*}$, defined as $w_{c}^{*}=\arg\min_{\hat{w}_{c}}\left\|\hat{J}(t)-r(t)-\alpha\hat{J}(t+1)\right\|$, and $w_{a}^{*}=\arg\min_{\hat{w}_{a}}\left\|r(t)+\alpha\hat{J}(t+1)\right\|$ if assuming $U_{c}=0$. The estimation errors of the critic and actor weights are defined by $\tilde{w}(t):=\hat{w}(t)-w^{*}$. Then Equations 20, 23, 26,31 define a dynamical system of estimation errors for a nonlinear function $F$:
$\tilde{w}(t+1)=\tilde{w}(t)-F\!\left(\hat{w}(t),\phi(t),\phi(t+1)\right)$ (32)
Definition 1.
A dynamical system is said to be uniformly ultimately bounded (UUB) with ultimate bound $b>0$, if for any $a>0$ and $t_{0}>0$, there exists a positive number $N=N(a,b)$ independent of $t_{0}$, such that $\|\tilde{w}(t)\|\leq b,\text{for all }t\geq N+t_{0}$, whenever $\|\tilde{w}(t_{0})\|\leq a$.
Theorem 1 (UUB of a discrete dynamical system).
If, for system (32), there exists a function $L(\tilde{w}(t),t)$ such that for all $\tilde{w}(t_{0})$ in a compact set $K$, $L(\tilde{w}(t),t)$ is positive definite and the first difference satisfies $\Delta L(\tilde{w}(t),t)<0,\text{for}\;\|\tilde{w}(t_{0})\|>b$, for some $b>0$ such that the $b$-neighborhood of $\tilde{w}(t)$ is contained in $K$, then the system is UUB and the norm of the state is bounded within a neighborhood of $b$.
V-B Network weights increments
The convergence of critic and actor weights is analyzed using gradient descent algorithm. The requirements on learning rate to stabilize the convergence are analyzed. The ultimate bound is also provided.
Assumption 1.
The sub-optimal weights are bounded as $\|w_{c}^{\ast}\|\leq w_{c}^{\max}$, $\|w_{a}^{\ast}\|\leq w_{a}^{\max}$.
Lemma 1.
Under Assumption 1, define $L_{1}(t)=\frac{1}{l_{c}}\operatorname{tr}\!\left[(\tilde{w}_{c}^{(2)}(t))^{T}\tilde{w}_{c}^{(2)}(t)\right]$, and its first difference $\Delta L_{1}(t)=L_{1}(t+1)-L_{1}(t)$ is
$\Delta L_{1}(t)$ $=l_{c}e_{c}^{2}(t)\|\phi_{c}(t)\|^{2}+\|e_{c}(t)-\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-\|\zeta_{c}(t)\|^{2}-e^{2}_{c}(t)$ (33)
where $\zeta_{c}(t)=\tilde{w}_{c}^{(2)}(t)\phi_{c}(t)$ is the approximation error of the critic output.
Proof.
According to (21), (22), (23) and the fact that $w_{c}^{\ast(2)}$ is time-invariant:
$\tilde{w}_{c}^{(2)}(t+1)$ $=\hat{w}_{c}^{(2)}(t+1)-w_{c}^{\ast(2)}=\tilde{w}_{c}^{(2)}(t)-l_{c}e_{c}(t)\phi^{T}_{c}(t)$ (34)
where $\phi_{c}(t)\in\mathbb{R}^{N_{h_{c}}\times 1}$, $e_{c}(t)$ is a scaler. The quadratic weight error is
$\text{tr}\left[\left(\tilde{w}_{c}^{(2)}(t+1)\right)^{T}\tilde{w}_{c}^{(2)}(t+1)\right]=$ $\text{tr}\left[\left(\tilde{w}_{c}^{(2)}(t)-l_{c}e_{c}(t)\phi^{T}_{c}(t)\right)^{T}\left(\tilde{w}_{c}^{(2)}(t)-l_{c}e_{c}(t)\phi^{T}_{c}(t)\right)\right]$ (35)
$=$ $\text{tr}\left[\left(\hat{w}_{c}^{(2)}(t)\right)^{T}\tilde{w}_{c}^{(2)}(t)\right]+\text{tr}\left[l^{2}_{c}e^{2}_{c}(t)\phi_{c}(t)\phi^{T}_{c}(t)\right]$
$\quad-l_{c}e_{c}(t)\text{tr}\left[\left(\tilde{w}_{c}^{(2)}(t)\right)^{T}\phi^{T}_{c}(t)\right]-l_{c}e_{c}(t)\text{tr}\left[\phi_{c}(t)\tilde{w}_{c}^{(2)}(t)\right]$
$=$ $\text{tr}\left[\left(\hat{w}_{c}^{(2)}(t)\right)^{T}\tilde{w}_{c}^{(2)}(t)\right]+\text{tr}\left[l^{2}_{c}e^{2}_{c}(t)\phi_{c}(t)\phi^{T}_{c}(t)\right]$
$\quad-2l_{c}e_{c}(t)\text{tr}\left[\phi_{c}(t)\tilde{w}_{c}^{(2)}(t)\right]$
$=$ $\text{tr}\left[\left(\hat{w}_{c}^{(2)}(t)\right)^{T}\tilde{w}_{c}^{(2)}(t)\right]+l^{2}_{c}e^{2}_{c}(t)\|\phi_{c}(t)\|^{2}$
$\quad-2l_{c}e_{c}(t)\tilde{w}_{c}^{(2)}(t)\phi_{c}(t)$
Since $\tilde{w}_{c}^{(2)}(t)\phi_{c}(t)$ is a scalar, the cross term becomes
$-2l_{c}e_{c}(t)\tilde{w}_{c}^{(2)}(t)\phi_{c}(t)=l_{c}\Bigl(\|e_{c}(t)-\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-\|\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-e_{c}^{2}(t)\Bigr)$ (36)
The difference $\Delta L_{1}(t)$ yields
$\Delta L_{1}(t)$ $=\frac{1}{l_{c}}\Bigl(\operatorname{tr}[(\tilde{w}_{c}^{(2)}(t+1))^{T}\tilde{w}_{c}^{(2)}(t+1)]-\operatorname{tr}[(\tilde{w}_{c}^{(2)}(t))^{T}\tilde{w}_{c}^{(2)}(t)]\Bigr)$ (37)
$=l_{c}e_{c}^{2}(t)\|\phi_{c}(t)\|^{2}+\|e_{c}(t)-\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-\|\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-e^{2}_{c}(t)$
$=\underbrace{l_{c}e_{c}^{2}(t)\|\phi_{c}(t)\|^{2}+\|e_{c}(t)-\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-\|\zeta_{c}(t)\|^{2}-e^{2}_{c}(t)}_{\Delta L_{1}(t)}$
where $\zeta_{c}(t)=\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)$. ∎
Lemma 2.
Under Assumption 1, define $L_{2}(t)=\frac{1}{l_{a}\gamma_{1}}\operatorname{tr}\!\left[(\tilde{w}_{a}^{(2)}(t))^{T}\tilde{w}_{a}^{(2)}(t)\right]$, and its first difference $\Delta L_{2}(t)=L_{1}(t+1)-L_{1}(t)$ is bounded by
$\Delta L_{2}(t)\leq\frac{1}{\gamma_{1}}\left(l_{a}e^{2}_{a}(t)\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}+4\|r(t)\|^{2}+8\alpha^{2}\left\|\zeta_{c}(t+1)\right\|^{2}+8\alpha^{2}\left\|w^{*(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\|C(t)\|^{2}\|\zeta_{a}(t)\|^{2}-e^{2}_{a}\right)$ (38)
where $\zeta_{a}(t)=\tilde{w}_{a}^{(2)}(t)\phi_{a}(t)$.
Proof.
According to (28), (29), (31) and the fact that $w^{*(2)}_{a}$ is time-invariant:
$\tilde{w}^{(2)}_{a}(t+1)$ $=\hat{w}^{(2)}_{a}(t+1)-w^{*(2)}_{a}=\tilde{w}^{(2)}_{a}(t)-l_{a}e_{a}(t)C(t)\phi^{T}_{a}(t)$ (39)
where $\phi_{a}\in\mathbb{R}^{N_{a_{h}}\times 1}$, $C(t)\in\mathbb{R}^{n\times 1}$ and its element $C_{k}(t)=\left(2Ru(t)+\left[\alpha\hat{w}^{(2)}_{c}(t)\phi^{\prime}(x(t+1))\hat{w}^{(1)}_{c}(t)\right]^{T}G(t-1)\right)_{k}$.
The trace of the quadratic weight error is
$\text{tr}\left[\left(\tilde{w}^{(2)}_{a}(t+1)\right)^{T}\tilde{w}^{(2)}_{a}(t+1)\right]$ $=\text{tr}\left[\left(\tilde{w}^{(2)}_{a}(t)-l_{a}e_{a}(t)C(t)\phi^{T}_{a}(t)\right)^{T}\left(\tilde{w}^{(2)}_{a}(t)-l_{a}e_{a}(t)C(t)\phi^{T}_{a}(t)\right)\right]$ (40)
$=\text{tr}\left[\left(\tilde{w}^{(2)}_{a}(t)\right)^{T}\tilde{w}^{(2)}_{a}(t)\right]+\text{tr}\left[l^{2}_{a}e^{2}_{a}(t)\phi_{a}(t)C^{T}(t)C(t)\phi^{T}_{a}(t)\right]$
$\qquad-l_{a}e_{a}(t)\text{tr}\left[\left(\tilde{w}^{(2)}_{a}(t)\right)^{T}C(t)\phi^{T}_{a}(t)\right]-l_{a}e_{a}(t)\text{tr}\left[\phi_{a}(t)C^{T}(t)\tilde{w}^{(2)}_{a}(t)\right]$
$=\text{tr}\left[\left(\tilde{w}^{(2)}_{a}(t)\right)^{T}\tilde{w}^{(2)}_{a}(t)\right]+\text{tr}\left[l^{2}_{a}e^{2}_{a}(t)\phi_{a}(t)C^{T}(t)C(t)\phi^{T}_{a}(t)\right]$
$\qquad-2l_{a}e_{a}(t)\text{tr}\left[\phi_{a}(t)C^{T}(t)\tilde{w}^{(2)}_{a}(t)\right]$
Consider the cross term, and use $\text{tr}(ABC)=\text{tr}(CAB)$:
$-2l_{a}e_{a}(t)\text{tr}\left(\phi_{a}(t)C^{T}(t)\tilde{w}^{(2)}_{a}(t)\right)$ $=-2l_{a}e_{a}(t)\text{tr}\left(C^{T}(t)\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)\right)$ (41)
$=-2l_{a}e_{a}(t)C^{T}(t)\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)$
$\leq l_{a}\left(\left\|e_{a}(t)-C^{T}(t)\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)\right\|^{2}-\left\|C^{T}(t)\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)\right\|^{2}-e^{2}_{a}(t)\right)$
$\leq l_{a}\left(\left\|e_{a}(t)-C^{T}(t)\zeta_{a}(t)\right\|^{2}-\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}-e^{2}_{a}(t)\right)$
where $\zeta_{a}(t)=\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)$.
Notice that $(a-b)^{2}-b^{2}\leq 2a^{2}+b^{2}$, and use $\|C^{T}(t)\zeta_{a}(t)\|^{2}\leq\|C^{T}(t)\|^{2}\|\zeta_{a}(t)\|^{2}$ by Cauchy–Schwarz inequality:
$\left\|e_{a}(t)-C^{T}(t)\zeta_{a}(t)\right\|^{2}-\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$ (42)
$\quad\leq 2\left\|e_{a}(t)\right\|^{2}+\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$
$\quad\leq 2\left\|r(t)+\alpha\hat{J}(t+1)\right\|^{2}+\|C^{T}(t)\zeta_{a}(t)\|^{2}$
$\quad\leq 2\left(\left\|r(t)\right\|+\alpha\left\|\hat{J}(t+1)\right\|\right)^{2}+\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$
$\quad\leq 4\|r(t)\|^{2}+4\alpha^{2}\left\|\hat{w}^{(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$
$\quad\leq 4\|r(t)\|^{2}+4\alpha^{2}\left\|(\tilde{w}^{(2)}_{c}+w^{*(2)}_{c})\phi_{c}(t+1)\right\|^{2}+\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$
$\quad\leq 4\|r(t)\|^{2}+8\alpha^{2}\left\|\tilde{w}^{(2)}_{c}\phi_{c}(t+1)\right\|^{2}+8\alpha^{2}\left\|w^{*(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$
$\quad\leq 4\|r(t)\|^{2}+8\alpha^{2}\left\|\zeta_{c}(t+1)\right\|^{2}+8\alpha^{2}\left\|w^{*(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$
$\quad\leq 4\|r(t)\|^{2}+8\alpha^{2}\left\|\zeta_{c}(t+1)\right\|^{2}+8\alpha^{2}\left\|w^{*(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\left\|C(t)\|^{2}\|\zeta_{a}(t)\right\|^{2}$
where $\zeta_{c}(t+1)=\tilde{w}^{(2)}_{c}\phi_{c}(t+1)$.
Therefore, the first difference is
$\Delta L_{2}(t)$ $=\frac{1}{l_{a}\gamma_{1}}\mathrm{tr}\!\left(\tilde{w}^{(2)}_{a}(t+1)^{T}\tilde{w}^{(2)}_{a}(t+1)-\tilde{w}^{(2)}_{a}(t)^{T}\tilde{w}^{(2)}_{a}(t)\right)$ (43)
$\leq\frac{1}{\gamma_{1}}\left(l_{a}e^{2}_{a}(t)\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}+4\|r(t)\|^{2}+8\alpha^{2}\left\|\zeta_{c}(t+1)\right\|^{2}+8\alpha^{2}\left\|w^{*(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\|C(t)\|^{2}\|\zeta_{a}(t)\|^{2}-e^{2}_{a}\right)$
Remark 2.
Note that $\Delta L_{2}$ depends on $C(t)$, which incorporates the control-effectiveness matrix $G_{t-1}$. Thus, a system with high control effectiveness, i.e. a more “sensitive” system, may exhibit a larger bound. Intuitively, this matches our expectation.
Remark 3.
$C(t)$ is defined differently from [Ref] given the objective function to improve the policy, i.e., $r(t)+\hat{J}(t+1)$ in IHDP and $\hat{J}(t)$ in ADHDP.
Remark 4.
If we introduce the normalization $\|(\hat{w}_{c}^{(2)}(t))^{T}C(t)\|^{2}=1$ and fix the input-layer weights, then Lemmas 1–2 yield the results of Liu et al. (2012).
Lemma 3.
Under Assumption 1, define $L_{3}(t)=\frac{1}{l_{c}\gamma_{2}}\text{tr}\left[(\tilde{w}_{c}^{(1)}(t))^{T}\,\tilde{w}_{c}^{(1)}(t)\right]$, and its difference $\Delta L_{3}(t)=L_{3}(t+1)-L_{3}(t)$ is bounded by
$\Delta L_{3}(t)\leq\frac{1}{\gamma_{2}}\Big(l_{c}e^{2}_{c}(t)\|a(t)\|^{2}\|y(t)\|^{2}+\|a(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+e^{2}_{c}(t)\Big)$ (44)
where $\gamma_{2}>0$ is a weighting factor and $a(t)\in\mathbb{R}^{N_{h_{c}}\times 1}$ is a vector with $a_{j}(t)=\hat{w}^{(2)}_{c,j}(t)\phi^{\prime}_{j}(t)$.
Proof.
Consider the update rule (20), the weight error is
$\tilde{w}^{(1)}_{c}(t+1)=\hat{w}^{(1)}_{c}(t+1)-w_{c}^{*(1)}=\tilde{w}^{(1)}_{c}(t)-l_{c}e_{c}(t)B(t)$ (45)
where $B(t)\in\mathbb{R}^{N_{h_{c}}\times m}$ and its element $B_{ji}(t)=\hat{w}^{(2)}_{c,j}(t)\phi_{j}^{\prime}(t)y_{i}(t)$, then $B(t)=a(t)y^{T}(t)$.
The quadratic weight error is
$\text{tr}\left[\tilde{w}^{(1)}_{c}(t+1)^{T}\tilde{w}^{(1)}_{c}(t+1)\right]$ $=\text{tr}\left[\left(\tilde{w}^{(1)}_{c}(t)-l_{c}e_{c}(t)B(t)\right)^{T}\left(\tilde{w}^{(1)}_{c}(t)-l_{c}e_{c}(t)B(t)\right)\right]$ (46)
$=\text{tr}\left[(\tilde{w}^{(1)}_{c}(t))^{T}\tilde{w}^{(1)}_{c}(t)\right]+\text{tr}\left[l_{c}^{2}e^{2}_{c}(t)B^{T}(t)B(t)\right]$
$\qquad-l_{c}e_{c}(t)\left[\left(\tilde{w}^{(1)}_{c}(t)\right)^{T}B(t)\right]-l_{c}e_{c}(t)\text{tr}\left[B^{T}(t)\tilde{w}^{(1)}_{c}(t)\right]$
$=\text{tr}\left[(\tilde{w}^{(1)}_{c}(t))^{T}\tilde{w}^{(1)}_{c}(t)\right]+\text{tr}\left[l_{c}^{2}e^{2}_{c}(t)B^{T}(t)B(t)\right]$
$\qquad-2l_{c}e_{c}(t)\text{tr}\left[B^{T}(t)\tilde{w}^{(1)}_{c}(t)\right]$
$=\text{tr}\left[(\tilde{w}^{(1)}_{c}(t))^{T}\tilde{w}^{(1)}_{c}(t)\right]+\text{tr}\left[l_{c}^{2}e^{2}_{c}(t)y(t)a^{T}(t)a(t)y^{T}(t)\right]$
$\qquad-2l_{c}e_{c}(t)\text{tr}\left[B^{T}(t)\tilde{w}^{(1)}_{c}(t)\right]$
$=\text{tr}\left[(\tilde{w}^{(1)}_{c}(t))^{T}\tilde{w}^{(1)}_{c}(t)\right]+l_{c}^{2}e^{2}_{c}(t)\|a(t)\|^{2}\|y(t)\|^{2}$
$\qquad-2l_{c}e_{c}(t)\text{tr}\left[B^{T}(t)\tilde{w}^{(1)}_{c}(t)\right]$
where by $\text{tr}(ABC)=\text{tr}(BCA)$:
$\text{tr}\left[B^{T}(t)\tilde{w}^{(1)}_{c}(t)\right]=$ $\text{tr}\left[y(t)a^{T}(t)\tilde{w}^{(1)}_{c}(t)\right]$ (47)
$=$ $\text{tr}\left[a^{T}(t)\tilde{w}^{(1)}_{c}(t)y(t)\right]$
$=$ $a^{T}(t)\tilde{w}^{(1)}_{c}(t)y(t)$
Use inequality $2ab\leq a^{2}+b^{2}$, the cross term becomes
$-2l_{c}e_{c}(t)a^{T}(t)\tilde{w}^{(1)}_{c}(t)y(t)$ $\leq l_{c}\left(\|a^{T}(t)\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+e^{2}_{c}(t)\right)$ (48)
$\leq l_{c}\left(\|a^{T}(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+e^{2}_{c}(t)\right)$
$\leq l_{c}\left(\|a(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+e^{2}_{c}(t)\right)$
Therefore, the first difference can be bounded by
$\Delta L_{3}(t)\leq\underbrace{\frac{1}{\gamma_{2}}\Big(l_{c}e^{2}_{c}(t)\|a(t)\|^{2}\|y(t)\|^{2}+\|a(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+e^{2}_{c}(t)\Big)}_{\Delta L_{3}(t)}$ (49)
Lemma 4.
Under Assumption 1, define $L_{4}(t)=\frac{1}{l_{a}\gamma_{3}}\operatorname{tr}\!\left[\tilde{w}^{(1)}_{a}(t)^{T}\tilde{w}^{(1)}_{a}(t)\right]$, and its first difference is bounded by
$\Delta L_{4}(t)\leq\frac{1}{\gamma_{3}}\Big(l_{a}e^{2}_{a}(t)\left\|C^{T}(t)D(t)\right\|^{2}\|x(t)\|^{2}+e^{2}_{a}(t)+\left\|C^{T}(t)D(t)\right\|^{2}\left\|\tilde{w}^{(1)}_{a}(t)\,x(t)\right\|^{2}\Big)$ (50)
where $\gamma_{3}>0$ is a weighting factor, and $D(t)\in\mathbb{R}^{2\times N_{h_{a}}}$, and its element $D_{kj}(t)=\hat{w}^{(2)}_{a,kj}(t)\sigma^{\prime}_{j}(x(t))$.
Proof.
According to (26), the weight errors are
$\tilde{w}^{(1)}_{a}(t+1)$ $=\hat{w}^{(1)}_{a}(t+1)-w_{a}^{*(1)}$ (51)
$=\tilde{w}^{(1)}_{a}(t)-l_{a}e_{a}(t)D^{T}(t)C(t)x^{T}(t)$
and
$\text{tr}\left[\left(\tilde{w}^{(1)}_{a}(t+1)\right)^{T}\tilde{w}^{(1)}_{a}(t+1)\right]$ $=\text{tr}\left[\left(\tilde{w}^{(1)}_{a}(t)-l_{a}e_{a}(t)D^{T}(t)C(t)x^{T}(t)\right)^{T}\left(\tilde{w}^{(1)}_{a}(t)-l_{a}e_{a}(t)D^{T}(t)C(t)x^{T}(t)\right)\right]$ (52)
$=\text{tr}\left[\left(\tilde{w}^{(1)}_{a}(t)\right)^{T}\tilde{w}^{(1)}_{a}(t)\right]+\text{tr}\left[l^{2}_{a}e^{2}_{a}(t)x(t)C^{T}(t)D(t)D^{T}(t)C(t)x^{T}(t)\right]$
$\qquad-l_{a}e_{a}(t)\text{tr}\left[\left(\tilde{w}^{(1)}_{a}(t)\right)^{T}D^{T}(t)C(t)x^{T}(t)\right]-l_{a}e_{a}(t)\text{tr}\left[x(t)C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)\right]$
$=\text{tr}\left[\left(\tilde{w}^{(1)}_{a}(t)\right)^{T}\tilde{w}^{(1)}_{a}(t)\right]+l^{2}_{a}e^{2}_{a}(t)\text{tr}\left[x(t)C^{T}(t)D(t)D^{T}(t)C(t)x^{T}(t)\right]$
$\qquad-2l_{a}e_{a}(t)\text{tr}\left[x(t)C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)\right]$
$=\text{tr}\left[\left(\tilde{w}^{(1)}_{a}(t)\right)^{T}\tilde{w}^{(1)}_{a}(t)\right]+l^{2}_{a}e^{2}_{a}(t)\|C^{T}(t)D(t)\|^{2}\|x(t)\|^{2}$
$\qquad-2l_{a}e_{a}(t)\text{tr}\left[x(t)C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)\right]$
Use $\text{tr}(ABC)=\text{tr}(BCA)$ and the fact that $C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)$ is a scaler:
$\text{tr}\left[x(t)C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)\right]$ $=\text{tr}\left[C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)\right]$ (53)
$=C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)$
The cross term can be transformed into the form using $2ab\leq a^{2}+b^{2}$ and Cauchy–Schwarz inequality:
$-2l_{a}e_{a}(t)C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)$ $\leq l_{a}\left(e^{2}_{a}(t)+\left\|C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)\right\|^{2}\right)$ (54)
$\leq l_{a}\left(e^{2}_{a}(t)+\|C^{T}(t)D(t)\|^{2}\|\tilde{w}^{(1)}_{a}(t)x(t)\|^{2}\right)$
Finally, we can obtain the upper bound
$\Delta L_{4}(t)\leq\underbrace{\frac{1}{\gamma_{3}}\Big(l_{a}e^{2}_{a}(t)\left\|C^{T}(t)D(t)\right\|^{2}\|x(t)\|^{2}+e^{2}_{a}(t)+\left\|C^{T}(t)D(t)\right\|^{2}\left\|\tilde{w}^{(1)}_{a}(t)\,x(t)\right\|^{2}\Big)}_{\Delta L_{4}(t)}$ (55)
V-C Convergence of the weights errors
We introduce the Lyapunov candidate
$L(t)=L_{1}(t)+L_{2}(t)+L_{3}(t)+L_{4}(t)$ (56)
Theorem 2 (Main Theorem).
Let the weights of the critic and actor be updated by gradient descent algorithm, and assume that the reward signal is a bounded semidefinite function. Under Assumption 1, the estimation errors $w_{a}^{\ast}-\hat{w}_{a}(t)$ and $w_{c}^{\ast}-\hat{w}_{c}(t)$ are uniformly ultimately bounded (UUB) if
$l_{c}<\min_{t}\frac{\gamma_{2}-1}{\gamma_{2}\big(\|\phi_{c}(t)\|^{2}+\frac{1}{\gamma_{2}}\|a(t)\|^{2}\|y(t)\|^{2}\big)}$ (57)
$l_{a}<\min_{t}\frac{\gamma_{3}-\gamma_{1}}{\gamma_{3}\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}+\gamma_{1}\|D(t)C^{T}\|^{2}\|x(t)\|^{2}}$ (58)
Proof of Theorem 2.
Collect all terms of $\Delta L(t)$ based on the results of Lemmas 1–4, $\Delta L(t)$ is bounded by
$\Delta L(t)\leq$ $(\Delta L_{1})=l_{c}\underbrace{e_{c}^{2}(t)}_{2}\|\phi_{c}(t)\|^{2}+\|e_{c}(t)-\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-\underbrace{\|\zeta_{c}(t)\|^{2}}_{1}-\underbrace{e^{2}_{c}(t)}_{2}$ (59)
$(\Delta L_{2})+\frac{1}{\gamma_{1}}\left(l_{a}\underbrace{e^{2}_{a}(t)}_{3}\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}+4\|r(t)\|^{2}+8\alpha^{2}\underbrace{\left\|\zeta_{c}(t+1)\right\|^{2}}_{1}+8\alpha^{2}\left\|w^{*(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\|C(t)\|^{2}\|\zeta_{a}(t)\|^{2}-\underbrace{e^{2}_{a}}_{3}\right)$ (60)
$(\Delta L_{3})+\frac{1}{\gamma_{2}}\Big(l_{c}\underbrace{e^{2}_{c}(t)}_{2}\|a(t)\|^{2}\|y(t)\|^{2}+\|a(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+\underbrace{e^{2}_{c}(t)}_{2}\Big)$ (61)
$\Delta L_{4}(t)+\frac{1}{\gamma_{3}}\Big(l_{a}\underbrace{e^{2}_{a}(t)}_{3}\left\|C^{T}(t)D(t)\right\|^{2}\|x(t)\|^{2}+\underbrace{e^{2}_{a}(t)}_{3}+\left\|C^{T}(t)D(t)\right\|^{2}\left\|\tilde{w}^{(1)}_{a}(t)\,x(t)\right\|^{2}\Big)$ (62)
After combining like terms, the expression simplifies to
$\Delta L(t)\leq$ $\underbrace{-\Big(\|\zeta_{c}(t)\|^{2}-\frac{8\alpha^{2}}{\gamma_{1}}\|\zeta_{c}(t+1)\|^{2}\Big)}_{1}-\underbrace{\Big(1-l_{c}\|\phi_{c}(t)\|^{2}-\frac{l_{c}}{\gamma_{2}}\|a(t)\|^{2}\|y(t)\|^{2}-\frac{1}{\gamma_{2}}\Big)e^{2}_{c}(t)}_{2}$ (63)
$-\underbrace{\Big(\frac{1}{\gamma_{1}}-\frac{l_{a}}{\gamma_{1}}\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}-\frac{l_{a}}{\gamma_{3}}\|D(t)C^{T}(t)\|^{2}\|x(t)\|^{2}-\frac{1}{\gamma_{3}}\Big)e^{2}_{a}(t)}_{3}$
$+\big\|w^{*(2)}_{c}(t)\phi_{c}(t)-r(t)-\alpha\hat{w}^{(2)}_{c}(t)\phi(t+1)\|^{2}+4\|r(t)\|^{2}+\frac{8\alpha^{2}}{\gamma_{1}}\|w_{c}^{*(2)}\phi_{c}(t+1)\|^{2}+\frac{1}{\gamma_{1}}\|C(t)\|^{2}\|\zeta_{a}(t)\|^{2}$
$+\frac{1}{\gamma_{2}}\|a(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}$
$+\frac{1}{\gamma_{3}}\|C^{T}(t)D(t)\|^{2}\|\tilde{w}^{(1)}_{a}(t)x(t)\|^{2}$
To ensure the negativity of the term 1:
$\frac{\gamma_{1}}{8\alpha^{2}}$ $>\frac{\|\zeta_{c}(t+1)\|^{2}}{\|\zeta_{c}(t)\|^{2}}$ (64)
$=\frac{\|\tilde{w}^{(2)}(t)\|^{2}\Big(\|\phi(x(t))\|^{2}+\|L\Delta x(t)\|^{2}\Big)}{\|\tilde{w}^{(2)}(t)\|^{2}\|\phi(x(t))\|^{2}}$
$=\frac{\phi(x(t))\|^{2}+\|L\Delta x(t)\|^{2}}{\|\phi(x(t))\|^{2}}$
$=1+\frac{\|L\Delta x(t)\|^{2}}{\|\phi(x(t))\|^{2}}$
where $\Delta x(t)=x(t+1)-x(t)$. The term $\frac{\|L\Delta x(t)\|^{2}}{\|\phi(x(t))\|^{2}}$ is the effect of the state transition from $x(t),x(t+1)$ on input to hidden layer of the critic, which can be proved to be upper bounded if $x(t),w^{(1)}_{c}(t)$ are upper bounded:
$\frac{\|L\Delta x(t)\|^{2}}{\|\phi(x(t))\|^{2}}\leq$ $\frac{\|L\|^{2}\|\Delta x(t)\|^{2}}{\|\tanh(w^{(1)}_{c}(t)x(t))\|^{2}}$ (65)
$\leq$ $\frac{\|w^{(1)}_{c}(t)x(t)\|^{2}{}\|\Delta x(t)\|^{2}}{\|\tanh(w^{(1)}_{c}(t)x(t))\|^{2}}$
$\leq$ $\frac{\|\overline{w}^{(1)}_{c}(t)\overline{x}(t)\|^{2}{}\|\Delta\overline{x}(t)\|^{2}}{\|\tanh(\overline{w}^{(1)}_{c}(t)\overline{x}(t))\|^{2}}$
$=$ $\delta$
given the monotonicity of $y=\frac{|x|^{2}}{|\tanh(x)|^{2}}$. $\overline{\|\Delta x(t)\|}$ is the upper bound of $\|\Delta x(t)\|$ and determined by the system dynamics and the length of discrete time step.
To ensure negativity of the term 2:
$1-l_{c}\|\phi_{c}(t)\|^{2}-\frac{l_{c}}{\gamma_{2}}\|a(t)\|^{2}\|y(t)\|^{2}-\frac{1}{\gamma_{2}}>0$ (66)
Therefore,
$l_{c}<\min_{t}\frac{\gamma_{2}-1}{\gamma_{2}\big(\|\phi_{c}(t)\|^{2}+\frac{1}{\gamma_{2}}\|a(t)\|^{2}\|y(t)\|^{2}\big)}$ (67)
In particular, $\gamma_{2}>1$.
Similarly, to ensure negativity of the term 3:
$\frac{1}{\gamma_{1}}-\frac{l_{a}}{\gamma_{1}}\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}-\frac{l_{a}}{\gamma_{3}}\|D(t)C^{T}\|^{2}\|x(t)\|^{2}-\frac{1}{\gamma_{3}}>0$ (68)
Therefore,
$l_{a}<\min_{t}\frac{\gamma_{3}-\gamma_{1}}{\gamma_{3}\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}+\gamma_{1}\|D(t)C^{T}\|^{2}\|x(t)\|^{2}}$ (69)
In particular, $\gamma_{3}>\gamma_{1}$.
The fourth term can be bounded as
$\big\|w_{c}^{*(2)}\phi_{c}(t)-r(t)-\alpha\hat{w}^{(2)}_{c}(t)\phi_{c}(t+1)\big\|^{2}$ $\leq 4\|w_{c}^{*(2)}\phi_{c}(t)\|^{2}+4r^{2}(t)+2\alpha^{2}\|\hat{w}^{(2)}_{c}(t)\phi_{c}(t+1)\|^{2}$ (70)
Let $\overline{C},\overline{w}_{a1},\overline{w}_{a2},\overline{w}_{c1},\overline{\phi}_{a},\overline{y},\overline{x},\overline{a},\overline{D}$ be upper bounds of $C,w_{a1},w_{a2},w_{c1},\phi_{a},y,x,a,D$, and let $w_{c2}=\max\{w_{c}^{*(2)},\,w^{(M)}_{c2}\}$, where $w^{(M)}_{c2}$ is the upper bound of $\hat{w}^{(2)}_{c}(t)$. Finally, we obtain the following bound:
$\big\|w_{c}^{*(2)}\phi_{c}(t)-r(t)-\alpha\hat{w}^{(2)}_{c}(t)\phi_{c}(t+1)\big\|^{2}+4\|r(t)\|^{2}+\frac{8\alpha^{2}}{\gamma_{1}}\|w_{c}^{*(2)}(t)\phi_{c}(t+1)\|^{2}+\frac{1}{\gamma_{1}}\|C(t)\|^{2}\|\zeta_{a}(t)\|^{2}$ (71)
$\quad+\frac{1}{\gamma_{2}}\|a(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+\frac{1}{\gamma_{3}}\|C^{T}(t)D(t)\|^{2}\|\tilde{w}^{(1)}_{a}(t)x(t)\|^{2}$
$\leq\Big(\frac{8\alpha^{2}}{\gamma_{1}}+4+2\alpha^{2}\Big)(\overline{w}_{c2}\overline{\phi}_{c})^{2}+8\overline{r}^{2}+\frac{1}{\gamma_{1}}(\overline{w}_{c(2)}\overline{C}\overline{w}_{a2}\overline{\phi}_{a})^{2}+\frac{1}{\gamma_{2}}(\overline{w}_{c1}\overline{ya})^{2}+\frac{1}{\gamma_{3}}(\overline{w}^{(2)}_{c}\overline{C}\overline{D}\overline{w}_{a1}\overline{x})^{2}=M$
The conclusion is, if $\gamma_{1}>8\alpha^{2}(1+\frac{\|L\Delta x(t)\|^{2}}{\|\phi(x(t))\|^{2}))}{\|\zeta(t)\|^{2}}$ in the term 1, and $l_{c}$, $l_{a}$ satisfying (67), 69, and $\|\tilde{w}^{(2)}_{c}(t)\|^{2}>\frac{M}{\|\phi(x(t))\|^{2}-\frac{8\alpha^{2}}{\gamma_{1}}\|\phi(x(t+1))\|^{2}}$, we obtain $\Delta L(t)<0$. Based on Theorem 1, this implies that the system of estimation errors is ultimately uniformly bounded.
VI Flight Control Simulation
This section introduces an online reinforcement learning framework for angle-of-attack tracking control. A cascaded control structure is employed for separate control of angle-of-attack tracking and pitch rate tracking [20, 22]. The learning process is a sample-based implementation of the optimization-based tracking control design in Section LABEL:section_optimization_based_control. The learning algorithm uses Incremental model-based heuristic dynamic programming (IHDP).
VI-A Aerial Vehicle Model
The dynamical model of aerial vehicles is given as [32]
$\dot{\alpha}$ $=\left(\frac{fgQS}{WV}\right)\cos(\alpha)[\phi_{z}(\alpha)+b_{z}\delta]+q$ (72)
$\dot{q}$ $=\left(\frac{fQSd}{I_{yy}}\right)[\phi_{m}(\alpha)+b_{m}\delta]$
where $\alpha$ is the angle of attack, $q$ is the pitch rate, $\delta$ is the control surface deflection. The aerodynamic coefficients are approximately computed by $b_{z}=-0.034,b_{m}=-0.206$, and
[…]
These approximations of $b_{z},b_{m},\phi_{z}(\alpha),\phi_{m}(\alpha)$ hold for $\alpha$ in the range of $\pm 20$ degrees. The physical coefficients are provided in Table I. In addition, the actuator dynamics are considered as a first order model with the time constant $0.005\text{s}$. The rate limit is 600 deg/s, and a control surface deflection limit is $\pm 20$ degrees.
TABLE I: Physical parameters (adapted from [32])
Notations Definition Value
$g$ acceleration of gravity 9.815 m/s2
$W$ weight 204.3 kg
$V$ speed 947.715 m/s
[…]
$Q$ dynamic pressure 29969.861 kg/m2
$S$ reference area 0.041 m2
$d$ reference diameter 0.229 m
VI-B Tracking errors dynamics
Define the tracking errors as $e_{1}=\alpha-\alpha_{\text{ref}}$, $e_{2}=q-q_{\text{ref}}$, one has
$\qquad\qquad\dot{e}_{1}=$ $\dot{\alpha}-\dot{\alpha}_{\text{ref}}$ (74)
$=$ $f_{1}(x_{1})+x_{2}+d_{1}-\dot{\alpha}_{\text{ref}}$
$=$ $f_{1}(x_{1})+e_{2}+q_{\text{ref}}+d_{1}-\dot{\alpha}_{\text{ref}}$
$\dot{e}_{2}=$ $\dot{q}-\dot{q}_{\text{ref}}$ (75)
$=$ $f_{2}(x_{1})+g_{2}u-\dot{q}_{\text{ref}}$
where $\alpha_{\text{ref}}$, $q_{\text{ref}}$ are the angle-of-attack reference and the pitch rate reference. where $x_{1},x_{2}]^{T}=[\alpha,q]^{T}$, $u=\delta$ are state and control input. The nonlinear functions are defined as
$f_{1}(x_{1})$ $=\left(\frac{fgQS}{WV}\right)\cos(\alpha)\phi_{z}(\alpha),\;d_{1}(x_{1},u)=\left(\frac{fgQS}{WV}\right)\cos(\alpha)b_{z}$ (76)
$f_{2}(x_{1})$ $=\left(\frac{fQSd}{I_{yy}}\right)\phi_{m}(\alpha),\qquad\quad g_{2}(x_{1},\delta)=\left(\frac{fQSd}{I_{yy}}\right)b_{m}$
The model LABEL:single_input_single_output_system is a single-input-single-output (SISO) system, as it has a single input that controls the state $x_{2}$ and subsequently influence the state $x_{1}$ through near-integrator dynamics, which exhibits dynamical oscillations when rate control is loose. The additional aerodynamic dynamics $f_{1}(\cdot)$ also disturb the angle-of-attack tracking error and the resulting control actions.
The virtual control defined as $q_{\text{ref}}(e_{1},\alpha,\delta)$ is associated with $e_{1},\alpha,\delta$, aiming to provide error-feedback control for stabilizing $e_{1}$ and from the intuitive point of view, compensate for the dynamics associated with $\alpha$ and $\delta$ in $e_{1}$-subsystem. In backstepping this nominal model is used to design the virtual control and its adaptive variants to ensure asymptotic convergence of the tracking errors under model uncertainties. In model-free RL settings, the design of $q_{\text{ref}}(e_{\alpha},\alpha,\delta)$ with guaranteed performance is infeasible as the dynamics are unknown. It is usually parameterized using a neural network and trained with samples. The same analysis applies to the design of $u$ in $e_{2}$-subsystem. The control surface deflection $\delta=u(e_{2},q,\alpha)$ is a function of $e_{2}$, $q$ and $\alpha$, in order to stabilize $e_{2}$ and compensate for the dynamics associated with $q,\alpha$.
VI-C Outer-loop and inner-loop agent
VI-C1 Incremental model
The incremental model of $q$-dynamics is developed as
$\Delta q_{t+1}=F_{t-1}\Delta q_{t}+G_{t-1}\Delta\delta_{t}$ (77)
where the subscript $t$ denotes the discrete-time index. The state increments are defined as $\Delta q_{t+1}=q_{t+1}-q_{t}$, $\Delta q_{t}=q_{t}-q_{t-1}$, the control increment is defined as $\Delta\delta_{t}=\delta_{t}-\delta_{t-1}$. The first-order partial derivatives are $F_{t-1}=\frac{\partial f_{2}}{\partial\alpha}|_{\alpha_{t-1},\delta_{t-1},q_{t-1}},G_{t-1}=\frac{\partial g_{2}u}{\partial\delta_{t}}|_{\alpha_{t-1},\delta_{t-1},q_{t-1}}$. The derivation is seen in Appendix VIII-A. In practice, the model is identified using linear Recursive Least Squares algorithm (Appendix VIII-B).
Remark 5.
The incremental model of the $\alpha$-dynamics is not required during the training of the outer-loop actor, even though its task is to control angle-of-attack. This control effectiveness from $q$ to $\alpha$ is implicitly captured in the estimated value function $\hat{V}_{2}(\cdot)$ by taking $q$ as an input.
VI-C2 Networked control laws
The outer-loop control law and inner-loop control laws are parameterized using two neural networks as $q_{\text{ref}}\left(e_{1},\alpha,\delta)=W_{1}(e_{1},\alpha,\delta\right)$, $\delta(e_{2},q,\alpha)=W_{2}(e_{2},q,\alpha)$. The nested control law is
$\delta(e_{2},q,\alpha)=\underbrace{W_{2}(q-\underbrace{W_{1}(e_{1},\alpha,\delta)}_{\text{outer-loop actor}},q,\alpha)}_{\text{inner-loop actor}}$ (78)
where $W_{1}(\cdot;w^{A_{1}}),W_{2}(\cdot;w^{A_{2}})$ are neural networks with trainable parameter sets $w^{A_{1}},w^{A_{2}}$. The cascaded structure is illustrated in Figure 1, along with a comparison to the monolithic actor.
VI-C3 Outer-loop agent
The outer-loop agent aims to learn a sub-optimal policy for pitch rate reference required to track the angle-of-attack reference. The network structures are shown in Figure LABEL:Figure_outer-loop_agent_networks. The critic takes angle-of-attack, its tracking error and pitch rate as inputs. The output is the estimated state-value $\hat{V}_{1(t)}$. The absolute-value function $\text{abs}(\cdot)$ is applied at the output layer to ensure positive definiteness. The actor uses control surface deflection as an additional input to attenuate the non-minimum phase dynamics. Its output is constrained using a scaled $\tanh(\cdot)$ function to satisfy the limits on the pitch-rate reference.
The critic is updated through TD learning to minimize a loss function:
$L^{C_{1}}_{t}=\frac{1}{2}\delta_{1(t)}^{T}\delta_{1(t)}$ (79)
where $\delta_{1(t)}=c_{1(t)}+\gamma\hat{V}_{1,\text{target}(t+1)}-\hat{V}_{1(t)}$ is the TD error, $\hat{V}_{1,\text{target}}(\cdot)$ is a target critic used to delay weight updates and stabilize the learning process.
The gradient of $L^{C_{1}}_{t}$ with respect to critic weights is
$\frac{\partial L^{C_{1}}_{t}}{\partial w^{C_{1}}_{t}}$ $=\frac{\partial L^{C_{1}}_{t}}{\partial\delta_{1(t)}}\frac{\partial\delta_{1(t)}}{\partial w^{C_{1}}_{t}}=-\delta_{1(t)}\frac{\partial\hat{V}_{1(t)}}{\partial w^{C_{1}}_{t}}$ (80)
The critic is updated by stepping in opposite direction to the gradient of the loss with learning rate $\eta^{C_{1}}_{t}$:
$w^{C_{1}}_{t}=w^{C_{1}}_{t}-\eta^{C_{1}}_{t}\frac{\partial L^{C_{1}}_{t}}{\partial w^{C_{1}}_{t}}$ (81)
VI-C4 Inner-loop agent
The inner-loop agent aims to learn a sub-optimal policy for control-surface deflection required to track the pitch-rate reference. The non-minimum-phase dynamics associated with $\delta$ may unintentionally affect the $\alpha$-dynamics. This is a key motivation for adopting a cascaded control structure, which isolates this effect from the inner-loop agent’s training process. As a result, the outer-loop agent is prevented from attempting to control angle-of-attack through the control-surface deflection directly.
Figure LABEL:Figure_inner_loop_agent_networks shows the network structures as single-hidden-layer fully-connected multi-layer-perceptions (MLP). The inputs of the critic are the pitch rate and the its tracking error. The output is the estimated state-value $\hat{V}_{2(t)}$. An absolute-value function $\text{abs}(\cdot)$ is applied at the output layer to ensure the positive definiteness. The actor uses angle-of-attack as an additional input to attenuate dynamic coupling in the $q$-dynamics. The output is the control surface deflection. A scaled $\tanh(\cdot)$ is applied at the output layer to constrain the action within actuator limits. The target critic and target actor share the same structures.
The critic is updated through Temporal Difference (TD) learning to minimize a loss function:
$L^{C_{2}}_{t}=\frac{1}{2}\delta_{2(t)}^{T}\delta_{2(t)}$ (82)
where $\delta_{2(t)}=c_{2(t)}+\gamma\hat{V}^{\prime}_{2(t+1)}-\hat{V}_{2(t)}$ is the TD error, $\hat{V}^{\prime}_{2}(\cdot)$ denotes the target critic, which is updated as a delayed version of $\hat{V}_{2}(\cdot)$ to stabilize policy learning.
The gradient of the critic’s loss with respect to its parameter vector is
$\frac{\partial L^{C_{2}}_{t}}{\partial w^{C_{2}}_{t}}$ $=\frac{\partial L^{C_{2}}_{t}}{\partial\delta_{2(t)}}\frac{\partial\delta_{2(t)}}{\partial w^{C_{2}}_{t}}=\delta_{2(t)}-\frac{\partial\hat{V}_{2(t)}}{\partial w^{C_{2}}_{t}}$ (83)
The critic is updated by
$w^{C_{2}}_{t+1}=w^{C_{2}}_{t}-\eta^{C_{2}}_{t}\frac{\partial L^{C_{2}}_{t}}{\partial w^{C_{2}}_{t}}$ (84)
The target critic is updated by
$w^{C^{\prime}_{2}}_{t+1}=\tau w^{C_{2}}_{t}+(1-\tau)w^{C^{\prime}_{2}}_{t}$ (85)
VI-C5 Actor
The actor is trained to minimize the estimated state-value function:
$L^{A_{2}}_{t}=c_{2(t)}+\gamma\hat{V}^{\prime}_{2(t+1)}$ (86)
$\delta^{*}_{t}=\arg\min_{\delta_{t}}L^{A_{2}}_{t}$ (87)
The gradient of the actor’s loss with respect to its parameter vector is
$\frac{\partial L^{A_{2}}_{t}}{\partial w^{A_{2}}_{t}}$ $=\frac{\partial L^{A_{2}}_{t}}{\partial\delta_{t}}\frac{\partial\delta_{t}}{\partial w^{A_{2}}_{t}}$ (88)
$=\left(2b\delta_{t}+\frac{\partial\hat{V}^{\prime}_{2}(e_{t+1},q_{t+1},\alpha_{t+1})}{\partial q_{t+1}}\frac{\partial q_{t+1}}{\partial\delta_{t}}\right)\frac{\partial\delta_{t}}{\partial w^{A_{2}}_{t}}$
$=\left(2b\delta_{t}+\frac{\partial\hat{V}^{\prime}_{2}(e_{t+1},q_{t+1},\alpha_{t+1})}{\partial q_{t+1}}\hat{G}_{t-1}\right)\frac{\partial\delta_{t}}{\partial w^{A_{2}}_{t}}$
The actor is updated by
$w^{A_{2}}_{t+1}=w^{A_{2}}_{t}-\eta^{A_{2}}_{t}\frac{\partial L^{A_{2}}_{t}}{\partial w^{A_{2}}_{t}}$ (89)
VI-C6 Training
The forward propagation of action $q_{\text{ref}}$ goes through the inner-loop actor to generate an action $\delta$ that interacts with the model, seen as an additional control effectiveness. This implies that the outer-loop actor is equivalently interacting with an extended model that includes the inner-loop actor (see Figure 1). The training sample is given by $(e_{1(t)},\alpha_{t},q_{t},\delta_{t},q_{\text{ref}},e_{1(t+1)},\alpha_{t+1},q_{t+1})$. Therefore, the gradient of the outer-loop actor propagates from $\hat{V}^{\prime}_{1(t+1)}$ and back through the inner-loop actor. The training of the inner-loop actor is independent of the outer-loop actor, since the target critic $\hat{V}^{\prime}_{2(t+1)}$ propagates gradients only to the inner-loop actor. Therefore, the samples $e_{1(t)},e_{1(t+1)}$ are not used during training. The training sample is given by $(e_{2(t)},q_{t},\alpha_{t},\delta_{t},e_{2(t+1)},q_{t+1})$. In this setting, the environment is treated as only the $e_{2}$-dynamics.
Figure 1: Cascaded actor versus monolithic actor. The former uses an outer-loop actor to generate a pitch-rate reference, which is then passed to an inner-loop actor that produces the control-surface deflection, providing an explicit mechanism for learning and tracking the pitch-rate reference. The latter is a fully connected multilayer perceptron (MLP) that approximates an end-to-end tracking policy. The pitch-rate reference is not explicitly represented by the monolithic actor. This end-to-end structure is commonly used for the stabilization of inverted pendulum and Van der Pol’s oscillator, without additional angular dynamics.
VI-D Action-smoothing techniques
This subsection introduces policy regularization for temporal smoothness technique and a low-pass filter to smoothen control actions. The first technique regularizes the actor to generate slow-varying actions, while the second technique filters the pitch rate reference.
VI-D1 Policy regularization for temporal action smoothness
The temporal action smoothness measure the extent to which the output changes when the actor input changes:
$L^{\text{TS}}_{1(t)}=\big|q_{\text{ref}(t+1)}-q_{\text{ref}(t)}\big|$ (90)
$L^{\text{TS}}_{2(t)}=\big|\delta_{t+1}-\delta_{t}\big|$ (91)
where $L^{\text{TS}}_{1(t)},L^{\text{TS}}_{2(t)}$ are temporal smoothness (TS) losses for the outer-loop and inner-loop actors, respectively.
To reduce computation load in online learning, a single sample is used at each step of policy update to compute TS losses. This contrasts with using a minibatch in offline learning [29].
The multi-objective optimization frameworks are formulated as
$q^{*}_{\text{ref}(t)}=\arg\min_{v_{t}}\Big(c_{1(t)}+\gamma\hat{V}^{\prime}_{1(t+1)}+\lambda_{1}L_{1(t)}^{\text{TS}}\Big)$ (92)
$\delta^{*}_{t}=\arg\min_{\delta_{t}}\Big(c_{2(t)}+\gamma\hat{V}^{\prime}_{2(t+1)}+\lambda_{2}L_{2(t)}^{\text{TS}}\Big)$ (93)
where $\lambda_{1}\geq 0,\lambda_{2}\geq 0$ are weights of TS losses.
VI-D2 Low-pass filter
A second-order filter is used to smooth the pitch rate reference [30]:
$\dot{d}_{1}$ $=d_{2}$ (94)
$\dot{d}_{2}$ $=-2\zeta\omega^{2}_{n}d_{2}-\omega^{2}_{n}(d_{1}-q_{\text{ref}})$
where $d_{1},d_{2}$ are filtered signal and its differentiation, $\omega_{n}$ is the natural frequency, $\zeta$ is the damping factor.
Remark 6.
The training samples are collected under the Markov assumption for the environment, i.e., the one-step future state is determined solely by the current state. However, the states of the filter depend on the past two states due to its integral dynamics, which violates this assumption when a filter is used during the training of the outer-loop actor. As a result, the gradients from the value function $\hat{V}_{1}$ are delayed by two time steps due to the filter. This issue can be mitigated by including the filter states as part of the environment state; however, this increases computational cost because of the higher dimensionality of the system and network inputs. To avoid delaying the gradient, this paper excludes the filter from the backpropagation of the policy through $\hat{V}_{1}$ and applies it only during the forward propagation of the cascaded actor.
Figure 2: Command-filtered cascaded online learning flight control system.
VII Simulation
This section presents simulation results for cascaded online-learning flight control systems that incorporate action-smoothness conditioning techniques. The aerial vehicle dynamics are numerically simulated through the fourth-order Runge Kutta method. The actuator dynamics are modeled as a first-order system $\tau\dot{\delta}_{t}+\delta_{t}=\delta_{c(t)},\tau=0.005$. To constrain actions, the outer-loop and inner-loop actors are scaled to $20^{\circ},20^{\circ}]$ and $20^{\circ}/\text{s},20^{\circ}/\text{s}]$. The reference of angle-of-attack is set as $\alpha_{\text{ref}(t)}=10^{\circ}\sin(\frac{2\pi}{T}t),T=10\text{s}$. The network weights are randomly initialized using a uniform distribution $\mathcal{U}(-0.01,0.01)$, which ensures an initially low-gain control law and slow learning to guard against inaccurate policy gradients. The persistence excitation signal is generated by adding Gaussian noise $\mathcal{N}(0,0.04)$ to the control surface deflection. The hyperparameters of the RL agents are listed in Table II. The baseline control system does not use either of action-smoothing techniques. We consider the tracking control performance in aspects of tracking errors, action smoothness, saturation level of activation function, actor sensitivity and one-step costs.
TABLE II: Hyperparameters of RL agents
Parameter Outer-loop agent Inner-loop agent
critic learning rate $\eta^{C_{1}}$, $\eta^{C_{2}}$ 0.1 0.1
actor learning rate $\eta^{A_{1}}$, $\eta^{A_{2}}$ $5\times 10^{-7}$ $10^{-7}$
discount factor $\gamma$ 0.6 0.6
delay factor $\tau$ 1 1
policy iteration number at each time step 3 3
hidden layer size 7 7
critic hidden-layer activation function tanh tanh
critic output layer activation function abs abs
actor activation function tanh tanh
optimizer Adam Adam
control effort weights $a,b$ $5\times 10^{-6}$ $10^{-5}$
smoothness loss weight $\lambda_{1},\lambda_{2}$ $9.3\times 10^{-4}$ $10^{-5}$
loss thresholds for terminating critic training $2.5\times 10^{-4}$ $5\times 10^{-5}$
maximal steps of critic and actor updates 50 50
VII-A Temporal-smoothed control system
VII-A1 Tracking control
In Figure 3 the baseline control system exhibits satisfactory tracking performance during the first 10 seconds, but then experiences growing oscillations of the states and actions that degrade the convergence of the tracking error, ultimately leading to divergence. This issue is addressed in [Ref] by constraining the actions to the interval [-5, 5], a hard constraint that does not prevent bang–bang control. In contrast, the temporal-smoothed control system produces smoother actions within [-6, 6] and avoids both divergence and bang–bang behavior by penalizing temporal variations in the actions. Figure 4 shows that the critic parameters of both the two systems converge to their sub-optimum within 10s, but the former continues to oscillate in response to large tracking errors. The actor parameters in the temporal-smoothed system exhibit slightly slower growth thanks to the stabilized tracking errors.
Figure 3: Tracking control performance.
Figure 4: Critic and actor parameters.
The frequency-domain characteristics of actions are captured by using the Fast Fourier Transform (FFT) algorithm, that converts the discrete-time action sequences into their frequency-domain representations, as given by the Discrete Fourier Transform $X[k]=\sum_{n=0}^{N-1}x[n]\,e^{-j\frac{2\pi}{N}kn},k=0,1,\ldots,N-1$. Figures 5 and 6 show the time-domain histories and the spectra over the 0–100 Hz range. The signals of $q_{\text{ref}}$ within 100 Hz are effectively attenuated using TS losses, and their amplitude of $q_{\text{ref}}$ is significantly reduced without compromising tracking performance. Similar results are observed for control surface deflection. A reason for the oscillation of action $\delta$ is that a switching tracking error $e_{2}$ induced by the switching $q_{\text{ref}}$ propagates through the lower-level actor to generate $\delta$.
Figure 5: Pitch rate in selected intervals.
Figure 6: Control surface deflection in selected intervals.
VII-A2 Activation function
The activation function $\tanh(\cdot)$ is saturated if its output approaches the maximum or minimum value due to large inputs. In the saturated region of $\tanh(\cdot)$, its gradient vanishes and approaches zero (see Figure 7). The vanishing derivative leads to a vanishing policy gradient, resulting in slow learning of the control policy. Another effect is that the actor produces actions near their maximum or minimum limits in response to small tracking errors, resulting in an almost bang–bang control law that degrades tracking performance and system stability.
The saturation of the actor is the result of an aggressive process of policy learning. This occurs for two main reasons: (1) a large learning rate. A large learning rate causes the actor’s weights to be updated with large increments, pushing the activation function into its saturation zone, a process that is irreversible due to vanishing gradients. (2) Underemphasized control effort in the one-step cost function. An one-step cost function that underemphasizes the control effort does not discourage the actor to use large actions during learning. The vanishing gradients make these aggressive actions stay and persist to degrade the tracking performance even as the tracking error decreases. Between 10 and 20s, although $\tanh(\cdot)$ is not saturated, significant switching actions occur, indicating that switching actions can happen even within the normal operating region of $\tanh(\cdot)$. Moreover, the time-varying dynamics and reference cause switching tracking errors, which propagate to the pitch-rate reference and control surface deflections through the cascaded actor.
Figure 7 shows the saturation level of $\tanh(\cdot)$ in output layers of actors. The baseline control system eventually makes the inputs reach to intervals [-4,-2], [2,4] (the outer-loop actor) and intervals [-3,-2], [2,3] (the lower-level actor). In these intervals, the derivative is $\tanh^{\prime}(\cdot)\leq 0.1$ and attenuates the overall policy gradient. As a comparison, the temporal-smoothed control system avoids reaching to the saturation regions. The inputs are kept within ’unsaturated’ interval [-2,2] (the outer-loop actor) and interval [-0.5,0.5] (the lower-level actor). This demonstrates the effectiveness of TS losses in avoiding saturation of $\text{tanh}(\cdot)$.
Figure 7: Input of the activation function $\tanh(\cdot)$ in output layers of actors.
VII-A3 Action increment
The one-step increment of pitch rate reference is
$\Delta q_{\text{ref}(t+1)}=\left|q_{\text{ref}(t+1)}-q_{\text{ref}(t)}\right|$ (95)
where $q_{\text{ref}(t+1)}=W_{1}(e_{1(t+1)},\alpha_{t+1},\delta_{t+1}),q_{\text{ref}(t)}=W_{1}(e_{1(t)},\alpha_{t},\delta_{t})$ are actors’ outputs.
The one-step increment of control surface deflection is
$\Delta\delta_{t+1}=\left|\delta_{t+1}-\delta_{t}\right|$ (96)
where $\delta_{t+1}=W_{2}(e_{2(t+1)},q_{t+1},\alpha_{t+1}),\delta_{t}=W_{2}(e_{2(t)},q_{t},\alpha_{t})$ are actors’ outputs .
Figure 8 shows the baseline control system suffers from fast-switching actions. Using a TS loss reduces the amplitude of the action increment to $|\Delta q_{\text{ref}(t+1)}|\leq 0.5\ ^{\circ}\text{s}$ and decreases the amplitude of $|\Delta\delta(t+1)|$, resulting in less saturated control actions.
Figure 8: Increments of actions.
VII-A4 Actor sensitivity
The sensitivity of the actor indicates how much, and in what direction, the output changes when the input increases slightly [34]. An actor with high sensitivity may overreact to its inputs, state variations and the tracking error. This is also correlated to robustness of the actor against sensor noise and uncertainties.
The sensitivity of the inner-loop and outer-loop actors is measured by
$K_{1(t)}$ $=\frac{\partial q_{\text{ref}(t)}}{\partial e_{1(t)}}$ (97)
$K_{2(t)}$ $=\frac{\partial\delta(t)}{\partial e_{2(t)}}$ (98)
The relation between the pitch rate reference and its sensitivity measure is
$|q_{\text{ref}(t+1)}-q_{\text{ref}(t)}|$ (99)
$\approx$ $\left|K_{1(t)}(e_{1(t+1)}-e_{1(t)})\right|$
$=$ $\left|K_{1(t)}\right|\cdot\left|e_{1(t+1)}-e_{1(t)}\right|$
Similarly, the relation between the control surface deflection and its sensitivity measure is
$|\delta_{(t+1)}-\delta_{(t)}|$ (100)
$\approx$ $\left|K_{2(t)}(e_{2(t+1)}-e_{2(t)})\right|$
$=$ $\left|K_{2(t)}\right|\cdot\left|e_{2(t+1)}-e_{2(t)}\right|$
Equations 99, 100 show that the action increments can be reduced through reducing sensitivity.
In Figure 9, the fast-growing $K_{1},K_{2}$ of the baseline control system indicate the actors are becoming significantly sensitive to variants of tracking errors. The oscillations are the results of actor saturation. The temporal-smoothed control system shows reduced oscillations because the TS losses penalize the increments of actions, which affects the actor sensitivity, as shown in Equations 99, 100.
Figure 9: Sensitivity measures.
VII-A5 Challenges
Although regularizing the policy using TS losses helps to generate smoother actions, the tuning of $\lambda_{1},\lambda_{2}$ remains empirical. Large values slow down the process of policy learning, while small values degrade action smoothness effect. Moreover, Figure 10 shows that the TS losses do not eliminate low-amplitude, high-frequency oscillations in $\Delta q_{\text{ref}(t+1)}$ and $\Delta\delta_{\text{ref}(t+1)}$, since the penalization effects diminish when the increments are small (see Equations 90, 91). In this situation, a low-pass filter can be used, as it effectively removes high-frequency components from a frequency-domain perspective, thereby attenuating these action oscillations. Meanwhile, incorporating a low-pass filter into the flight control system does not interfere with the training process of actors.
VII-B Command-filtered and temporal-smoothed control system
We introduce a low-pass filter into the cascaded online learning flight control system, which has been demonstrated effective in filtering pitch rate reference in backstepping [5]. The natural frequency and damping ratio are $\omega_{n}=20\text{rad/s}$ and $\zeta=0.7$.
VII-B1 Tracking control
Figure 10 compares the temporal-smoothed control system with its counterpart that includes a filter. The amplitude of the control surface deflection oscillation exceeds that of the pitch-rate reference oscillation, due to the proportional amplification of $e_{2}$ by the inner-loop actor. The filter is capable of attenuating oscillations in pitch rate reference, leading to smooth $e_{2}$ and smooth control surface deflection through the inner-loop actor. Figure 11 presents detailed comparisons of $q_{\text{ref}}$ over selected time periods. The $q_{\text{ref}}$ signals in the 10-40Hz range are effectively attenuated by the filter. Figure 12 provides detailed comparisons on $\delta$ over selected time periods. The $\delta$ signals in the 10-40Hz range are effectively attenuated. These observations account for the chosen filter parameters. Figure 13 shows the saturation level of $\tanh(\cdot)$ in the output layer of actors. Both control systems prevent saturation of $\tanh(\cdot)$, ensuring that its derivative satisfies $\tanh^{\prime}(\cdot)\geq 0.4$ (the outer-loop actor) and $\tanh^{\prime}(\cdot)\geq 0.8$ (the inner-loop actor). The comparisons on $K_{1},K_{2}$ in Figure 14 show that applying a filter slows their growth and reduces oscillations, contributing to a more robust controller.
Figure 10: Tracking performance comparison.
Figure 11: Pitch rate in selected intervals.
Figure 12: Control surface deflection in selected intervals.
Figure 13: Input of the activation function $\tanh(\cdot)$ in the output layer of actors.
Figure 14: Sensitivity measures.
VII-C Terminate and restart learning
VII-C1 Termination
The previous analysis assumes continual learning; that is, the agents never terminate learning during system operation. The issue lies in that growing control gains can cause state oscillations, even when tracking errors are small. Therefore, it is intuitive to terminate and restart learning based on tracking performance over a recent time period. Specifically, the actor is fixed if the accumulated tracking error over a period $T_{s}$ remains within a specified threshold:
$\overline{e}_{s}\leq\epsilon$ (101)
where $\overline{e}_{s}=\frac{1}{n_{s}}\sum^{n_{s}-1}_{i=0}|e_{t-i}|$ is the average of absolute tracking errors, $n_{s}=T_{s}/\Delta T$ is the number of samples, $\epsilon>0$ is an error threshold. A large error threshold makes learning terminated early when tracking errors are still large. Define the online-learning state indicator as $\tau\in{1,0}$, with $\tau=1$ denoting that learning is off and $\tau=0$ denoting that learning is on.
During the initial phase, when $t<T_{\text{s}}$ and there are not enough samples to compute $\overline{e}_{\text{s}}$, learning remains active:
$\tau=\begin{aligned} &0,\;\text{if}\;\;t<T_{\text{s}}\end{aligned}$ (102)
Figure 15: Termination case (left) and restart case (right). The shaded area represents an average performance threshold. The actual average tracking errors must remain below this threshold to terminate learning, and exceed it to trigger a restart.
VII-C2 Restart
The learning must restart when the historical control performance degrades, typically occurring when the control system falls into faults and uncertainties. The restart condition determines in what case to restart learning. The evaluation period, denoted by $T_{r}$, should be carefully chosen to ensure it qualifies to represent the required control performance. A long evaluation period would delays the restart, preventing a timely response to performance degradation, while a short evaluation period makes the restart decision overly sensitive, potentially triggered by measurement noise or short-term peaks. As a result, the learning state may switch frequently between on and off.
The restart condition is
$\overline{e}_{r}\geq\sigma$ (103)
where $\overline{e}_{r}=\frac{1}{n_{r}}\sum^{n_{r}-1}_{i=0}|e_{t-i}|$ is the average of absolute tracking errors, $n_{r}=T_{r}/\Delta T$ is the number of samples, $\sigma>0$ is an error threshold. The sketch maps are provided in Figure 15. The pseudocode is seen in Algorithm 1.
The outer-loop and inner-loop agents have independent termination and restart conditions. Their parameters are seen in Table III. Figure 16 shows the histories of average tracking errors and the indicators. The outer-loop agent is turned off at 29s and restarted at 31.5s, while the inner-loop agent is turned off at 22s and is never restarted.
Data: $T_{s},\epsilon,T_{r},\sigma$, $t$, $\Delta T$
1 Initialization termination indicator $\tau_{-1}=0$
2 for each environment step $t\leftarrow$ to $t_{\text{max}}$ do
3 if $\tau_{t-1}=0$ then
4    if $t\leq T_{s}$ then
5       $\tau_{t}=0$
6    else
7       if $\overline{e}_{s}\leq\epsilon$ then
8          $\tau_{t}=1$
9       else
10          $\tau_{t}=0$
11         end if
12
13      end if
14
15   end if
16
17 end for
18if $\tau_{t-1}=1$ then
19 if $\overline{e}_{r}\geq\sigma$ then
20    $\tau_{t}=0$
21 else
22    $\tau_{t}=1$
23   end if
24
25 end if
Algorithm 1 Termination and restart algorithm
TABLE III: Parameters
Parameter Outer-loop agent Inner-loop agent
$T_{s}$ 10s 10s
$\epsilon$ 0.25 0.5deg/s
$T_{r}$ 0.2s 0.1s
$\sigma$ 0.8 2deg/s
Figure 16: Termination and restart measures.
VIII Conclusion
We introduced two action-smoothing techniques into the cascaded online learning flight control system: temporal policy regularization and low-pass filter. The temporal policy regularization is verified through simulation to reduce the amplitude of pitch rate reference and control surface deflection. The low-pass filter is verified capable of removing the high-frequency signals of pitch rate reference, which cannot be accomplished by only using temporal policy regularization. The combination of these two methods enables a sufficient-level action smoothness during online learning. These techniques are necessary for online learning-based control of aerial vehicle systems with time-varying dynamics, suffering from oscillations of states and actions caused by switching tracking error.
Appendix
VIII-A Derivation of incremental model
The continuous nonlinear system is given as
$\dot{x}(t)=f(x(t),u(t))$ (104)
where the dynamics $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is a smooth function associated with state vector $x\in\mathbb{R}^{n}$ and input vector $u\in\mathbb{R}^{m}$. $n,m$ are positive integers denoting the dimensions of the state and action spaces.
Taking Taylor expansion for $f(x(t),u(t))$ at $t_{0}$:
$f(x(t),u(t))=$ $f(x(t_{0}),u(t_{0}))+F(t_{0})(x(t)-x(t_{0}))+G(t_{0})(u(t)-u(t_{0}))$ (105)
$+O\left[(x(t)-x(t_{0}))^{2},(u(t)-u(t_{0}))^{2}\right]$
where $F(t_{0})=\partial f(x,u)/\partial x|_{x(t_{0}),u(t_{0})}\in\mathbb{R}^{n\times n}$, $G(t_{0})=\partial f(x,u)/\partial u|_{x(t_{0}),u(t_{0})}\in\mathbb{R}^{n\times m}$ are first-order partial derivatives at $t_{0}$. $O(\cdot)$ is the summed higher-order terms.
The summed higher-order terms $O(\cdot)$ have a relatively small effect on the change from $x_{t}$ to $x_{t+1}$ compared with the first-order terms. Moreover, these higher-order derivatives are difficult to identify from data, which poses challenges for practical application of an incremental model that includes them. For these two reasons, we omit $O(\cdot)$ when defining an effective and practical model for the nonlinear system:
$f(x(t),u(t))=$ $f(x(t_{0}),u(t_{0}))+F(t_{0})(x(t)-x(t_{0}))+G(t_{0})(u(t)-u(t_{0}))$ (106)
Using Equation 104:
$\dot{x}(t)=\dot{x}(t_{0})+F(t_{0})\left[x(t)-x(t_{0})\right]+G(t_{0})\left[u(t)-u(t_{0})\right]$ (107)
The time derivatives are approximated by $\dot{x}(t)\approx\frac{x(t+T)-x(t)}{T}$ and $\dot{x}(t_{0})\approx\frac{x(t_{0}+T)-x(t_{0})}{T}$, where $T$ is a constant time step. Then
$\frac{x(t+T)-x(t)}{T}\approx\frac{x(t_{0}+T)-x(t_{0})}{T}+F(t_{0})\left[x(t)-x(t_{0})\right]+G(t_{0})\left[u(t)-u(t_{0})\right]$ (108)
Define a discretization approach over $T$ by setting $t_{0}=(k-1)T$, $x(t_{0})=x((k-1)T)\triangleq x_{k-1}$. Let $x(t)$ denote the next-step state following $x(t_{0})$, then $x(t)=x(t_{0}+T)\triangleq x_{k}$.
A discretized version of system 108 is
$x_{k+1}-x_{k}\approx x_{k}-x_{k-1}+F_{k-1}(x_{k}-x_{k-1})T+G_{k-1}(u_{k}-u_{k-1})T$ (109)
Use denotions $\Delta x(k+1)=x(k+1)-x(k)$, $\Delta x(k)=x(k)-x(k-1)$, $\Delta u(k)=x(k)-x(k-1)$:
$\Delta x_{k+1}=\Delta x_{k}+F_{k-1}\Delta x_{k}T+G_{k-1}\Delta u_{k}T$ (110)
VIII-B Recursive Least Squares identification
The linear forgetting Recursive Least Squares (RLS) algorithm is applied to identify matrices $F_{t-1},G_{t-1}$ are identified online. Define the augmented state as $\mathbf{x}_{t}=\left[\Delta q_{t},\Delta\delta_{t}\right]^{T}$ and estimated parameter vector as $\hat{\theta}_{t-1}=[\hat{F}_{t-1},\hat{G}_{t-1}]^{T}$. The one-step state prediction is made as $\Delta\hat{\mathbf{x}}_{t+1}^{T}=\mathbf{x}_{t}^{T}\hat{\theta}_{t-1}$. The error between the true state and the predicted state is denoted by $\varepsilon_{t}$ and defined as $\varepsilon_{t}=\Delta\mathbf{x}_{t+1}^{T}-\Delta\hat{\mathbf{x}}_{t+1}^{T}$. The parameter matrix is updated by
$\hat{\theta}_{t}=\hat{\theta}_{t-1}+P_{t}\mathbf{x}_{t}\varepsilon_{t}$ (111)
$P^{-1}_{t}$ $=\alpha P^{-1}_{t-1}+\mathbf{x}_{t}\mathbf{x}^{T}_{t}$ (112)
$=\sum^{t}_{j=1}\alpha^{t-j}\mathbf{x}_{j}\mathbf{x}^{T}_{j}+\alpha^{t}P^{-1}_{0}$
$=\sum^{t}_{j=1}\alpha^{t-j}\mathbf{x}_{j}\mathbf{x}^{T}_{j}+I\alpha^{t}\lambda_{\text{min}}(0)$
The initial auxiliary matrix is set as $P_{0}=I/\lambda_{\text{min}}(0)$, where $0<\lambda_{\text{min}}(0)\ll 1$ is the minimum eigenvalue of $P^{-1}_{0}$. The initial parameter matrix is $\hat{\theta}_{0}$. The forgetting factor is $0\ll\alpha<1$.