### v5

# Temporally Smoothed Incremental Heuristic Dynamic Programming for Command-filtered Cascaded Online Learning Flight Control System

Abstract
Model-free reinforcement learning has demonstrated the capability to adjust control laws using data-constructed policy gradients. Its application to online flight control commonly adopts a cascaded control architecture, which enables efficient policy learning for angle-of-attack tracking using control-surface deflections. However, the tracking performance and system stability are degraded by oscillatory actions. To address this issue, we incorporate policy regularization for temporal action smoothness and a low-pass filter to attenuate both the amplitude and oscillatory frequency of actions. The Fast Fourier Transform (FFT) algorithm is used to analyze action smoothness in the frequency-domain representation. Flight control simulations demonstrate the effectiveness of the two proposed techniques.
Index Terms:
reinforcement learning, symmetry, data augmentation, approximate value iteration, flight control.
I Introduction
Model-free reinforcement learning (RL) has been applied to flight control law optimization without using an accurate system model for state prediction. This approach significantly reduce the effort required to model the complex dynamics of aerial vehicles in control design, as the policy gradient is constructed using measured data. Model-based RL implemented by ADP commonly uses a nonlinear approximator, commonly an artificial neural network (ANN) to approximate system dynamics, in order to make accurate state prediction to achieve dynamic programming. However, the ADP-based implementation of model-free RL does not require learning a full system model, since only the first-order derivatives of the system dynamics are needed to construct the policy gradients. This motivates the use of a simpler model with fewer parameters than a nonlinear approximator, while still capturing the first-order derivatives, such as an incremental model, thereby simplifying the training process. The incremental model is a local linear approximation of a nonlinear system derived using Taylor expansion. The combination with Recursive Least Squares (RLS) identification in ADP framework has led to the development of incremental model-based ADP (IADP) methods, enabling computationally efficient online policy learning. Compared with action dependent heuristic dynamic programming (ADHDP), which implicitly approximates the control effectiveness within a state-action value ($Q$) function, the RLS algorithm quickly and accurately identifies the local control effectiveness represented in an incremental model, leading to a more reliable construction of the policy gradient, as demonstrated in simple cases [Ref]. For example, incremental model-based heuristic dynamic programming (IHDP) requires only the signs of the control effectiveness matrix $G$ to characterize the gradient directions, thereby using less model information than other IADP variants, such as incremental model-based dual heuristic programming (IDHP) and incremental model-based global dual heuristic programming (IGDHP).
The theoretical analysis of online ADHDP has been studied in [Ref], where the gradient descent algorithm is used to update critic and actor over time. The weights are proved to be ultimately uniformly boundedness (UUB) to the sub-optimal values. learning process involves three time steps. IHDP updates the critic in the same manner as ADHDP, but updates the actor using an estimated value function reconstructed according to the Bellman equation, rather than a state–action ($Q$) function as in ADHDP. In this study, we formulate the update laws of IHDP over two time steps (which is three time steps in [Ref]) and analyze the UUB convergence of the estimated weights to the suboptimal values, taking into account temporal action smoothness as discussed later. As a result, the policy gradients are constructed from the Bellman error and the temporal smoothness loss.
IHDP have been applied to online tracking policy optimization for pitch rate, angle-of-attack (AOA), and altitude for the Flying-V, the Cessna Citation business jet, and the F-16 fighter model. For AOA tracking control, a cascaded control structure is commonly used to regulate angle-of-attack tracking in the outer-loop system and pitch-rate tracking in the inner-loop system, allowing a hierarchical control strategy that is able to explicitly incorporates pitch-rate constraints. As a comparison, the monolithic structure can also constrain the pitch rate through penalization in the reward design, but its does not learn a pitch rate reference, leading to degraded rate control performance. In addition, the rate control performance is sensitive to the weight of the pitch-rate loss term. Therefore, it is important to develop a well-designed pitch-rate reference for ensuring reliable control performance. An illustrative example is backstepping, where a stabilizing pitch-rate reference policy is derived from a nominal model which ensures the convergence of tracking errors, with accompanying adaptive laws to handle model uncertainties.
However, designing a stabilizing pitch-rate reference is challenging in the model-free IHDP framework, since it relies solely on measured data for pitch-rate reference optimization, which usually leads to a destabilizing reference during learning. Another challenge lies in the frequently switching tracking errors near zero caused by time-varying nonlinear dynamics and time-varying $\alpha$ reference signals, which is more serious than in stabilizing linear time-invariant (LTI) systems. These switching tracking errors, in turn, generate switching pitch-rate references through the outer-loop actor, further perturbing the control surface deflections through the inner-loop actor. This switching phenomenon is also observed in sliding mode control systems.
A third factor contributing to oscillations of flight control systems is actor saturation, in which case the actor uses maximum or minimum actions responding to switching tracking errors near zero, known as bang–bang control. Meanwhile, the saturated actor yields vanishing policy gradients due to the activation functions, causing the weights to become nearly fixed and difficult to adjust flexibly. To mitigate this problem, discounted learning rates have been used to slow down actor updates over time, reducing the likelihood of entering the saturation region. The disadvantages lie in that discounted learning rate is selected empirically and is sensitive to its effectiveness. Also, it does not completely prevent actor saturation as learning progresses. The other solution is to terminate actor updates by fixing the weights once the required control performance is achieved, either by setting a termination time or a tracking error threshold, a strategy that has not yet been studied.
In summary, these practical challenges compromise the tracking control performance and stability of the flight control system, giving rise to a critical issue concerning action smoothness during learning.
Temporal smoothness of a control policy has been studied in Conditioning for action policy smoothness (CAPS), an action-smoothening technique that regularizing action smoothness in temporal and spatial scales. The first application of temporal smoothness loss is on offline-learning quadrotor control [Ref], where the amplitude of switching motor actions is significantly reduced. The subsequent applications are for offline-learning of Flying-V, and Cessna Citation flight controllers. Unfortunately, the computation of a spatial smoothness loss requires generating spatially biased state samples using a prior state distribution, which is difficult for online learning with limited samples. Therefore, we consider only learning temporal-smooth flight control policy. The low-pass filter has been used to remove high-frequency noise in command signals in command-filtered backstepping, Incremental Nonlinear Dynamic Inversion (INDI). The advantage of a filter is smoothening high-frequency commands by tuning the natural frequency, even the amplitude of the command is small. This compensates the disadvantage of temporal smoothness in removing low-frequency commands.
The main contributions of this study are summarized as:
  • We develop temporally smoothed incremental heuristic dynamic programming (TS-IHDP), a model-free online learning method to learn an action-increment-constrained control policy. The UUB convergence of the estimated weights of the critic and actor to suboptimal values is proved via trace analysis.
  • The cascaded online learning flight control simulations demonstrate that TS-IHDP and command-filtered TS-IHDP generate smooth pitch rate references and control surface deflections, reducing system oscillations and improving the system stability.
  • Termination and restart conditions are designed to adaptively control the online learning process based on tracking errors over sliding past-time windows.
The remainder of this paper is structured as follows. Section II introduces the incremental model and recursive least square identification. Section III formulates the TS-IHDP method, with weight updates using gradient descent. The convergence of weights is proved in Section IV. Section V designs a cascaded online learning flight control system. Section VI presents simulation results. Section VII concludes this paper.
II Preliminaries
II-A Incremental Model
The continuous nonlinear system is given as
$\dot{x}(t)=f(x(t),u(t))$ (1)
where the dynamics $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is a smooth function associated with state vector $x\in\mathbb{R}^{n}$ and input vector $u\in\mathbb{R}^{m}$. $n,m$ are positive integers denoting the dimensions of the state and action spaces.
Taking Taylor expansion for $f(x(t),u(t))$ at $t_{0}$:
$f(x(t),u(t))=$ $f(x(t_{0}),u(t_{0}))+F(t_{0})(x(t)-x(t_{0}))+G(t_{0})(u(t)-u(t_{0}))$ (2)
$+O\left[(x(t)-x(t_{0}))^{2},(u(t)-u(t_{0}))^{2}\right]$
where $F(t_{0})=\partial f(x,u)/\partial x|_{x(t_{0}),u(t_{0})}\in\mathbb{R}^{n\times n}$, $G(t_{0})=\partial f(x,u)/\partial u|_{x(t_{0}),u(t_{0})}\in\mathbb{R}^{n\times m}$ are first-order partial derivatives at $t_{0}$. $O(\cdot)$ is the summed higher-order terms.
The summed higher-order terms $O(\cdot)$ have a relatively small effect on the change from $x_{t}$ to $x_{t+1}$ compared with the first-order terms. Moreover, these higher-order derivatives are difficult to identify from data, which poses challenges for practical application of an incremental model that includes them. For these two reasons, we omit $O(\cdot)$ when defining an effective and practical model for the nonlinear system:
$f(x(t),u(t))=$ $f(x(t_{0}),u(t_{0}))+F(t_{0})(x(t)-x(t_{0}))+G(t_{0})(u(t)-u(t_{0}))$ (3)
Using Equation 1:
$\dot{x}(t)=\dot{x}(t_{0})+F(t_{0})\left[x(t)-x(t_{0})\right]+G(t_{0})\left[u(t)-u(t_{0})\right]$ (4)
The time derivatives are approximated by $\dot{x}(t)\approx\frac{x(t+T)-x(t)}{T}$ and $\dot{x}(t_{0})\approx\frac{x(t_{0}+T)-x(t_{0})}{T}$, where $T$ is a constant time step. Then
$\frac{x(t+T)-x(t)}{T}\approx\frac{x(t_{0}+T)-x(t_{0})}{T}+F(t_{0})\left[x(t)-x(t_{0})\right]+G(t_{0})\left[u(t)-u(t_{0})\right]$ (5)
Define a discretization approach over $T$ by setting $t_{0}=(k-1)T$, $x(t_{0})=x((k-1)T)\triangleq x_{k-1}$. Let $x(t)$ denote the next-step state following $x(t_{0})$, then $x(t)=x(t_{0}+T)\triangleq x_{k}$.
A discretized version of system 5 is
$x_{k+1}-x_{k}\approx x_{k}-x_{k-1}+F_{k-1}(x_{k}-x_{k-1})T+G_{k-1}(u_{k}-u_{k-1})T$ (6)
Use denotions $\Delta x(k+1)=x(k+1)-x(k)$, $\Delta x(k)=x(k)-x(k-1)$, $\Delta u(k)=x(k)-x(k-1)$:
$\Delta x_{k+1}=\Delta x_{k}+F_{k-1}\Delta x_{k}T+G_{k-1}\Delta u_{k}T$ (7)
II-B Recursive Least Squares identification
The Recursive Least Squares (RLS) algorithm identifies matrices $F_{t-1},G_{t-1}$ online. Define the augmented state as $\mathbf{x}_{t}=\left[\Delta q_{t},\Delta\delta_{t}\right]^{T}$ and estimated parameter as $\hat{\theta}_{t-1}=[\hat{F}_{t-1},\hat{G}_{t-1}]^{T}$. The one-step state prediction is $\Delta\hat{\mathbf{x}}_{t+1}^{T}=\mathbf{x}_{t}^{T}\hat{\theta}_{t-1}$. The error between the true state and predicted state is $\varepsilon_{t}=\Delta\mathbf{x}_{t+1}^{T}-\Delta\hat{\mathbf{x}}_{t+1}^{T}$. The estimated parameter is updated by
$\hat{\theta}_{t}=\hat{\theta}_{t-1}+P_{t}\mathbf{x}_{t}\varepsilon_{t}$ (8)
$P^{-1}_{t}$ $=\alpha P^{-1}_{t-1}+\mathbf{x}_{t}\mathbf{x}^{T}_{t}$ (9)
$=\sum^{t}_{j=1}\alpha^{t-j}\mathbf{x}_{j}\mathbf{x}^{T}_{j}+\alpha^{t}P^{-1}_{0}$
$=\sum^{t}_{j=1}\alpha^{t-j}\mathbf{x}_{j}\mathbf{x}^{T}_{j}+I\alpha^{t}\lambda_{\text{min}}(0)$
The initial auxiliary matrix is set as $P_{0}=I/\lambda_{\text{min}}(0)$, where $0<\lambda_{\text{min}}(0)\ll 1$ is the minimum eigenvalue of $P^{-1}_{0}$. The initial parameter matrix is $\hat{\theta}_{0}$. The forgetting factor is $0\ll\alpha<1$.
III Temporally Smoothed Incremental Model-based Heuristic Dynamic Programming
III-A Problem formulation
Consider a discrete-time nonlinear dynamical system as
$x(t+1)=f\big(x(t),\,u(t)\big),$ (10)
where $x\in\mathbb{R}^{m}$ is the state vector and $u\in\mathbb{R}^{n}$ is the control vector.
The cost-to-go function is expressed as
$J(x(t))=\sum_{i=t}^{\infty}\alpha^{\,i-t}r\big(x(i),u(i)\big),$ (11)
where $\alpha\in(0,1]$ is the discount factor, $r(x(t),u(t))$ is the reward function. We require $r(t)=r(x(t),u(t))$ to be a bounded semidefinite function of the state and control so that the cost function is well defined. One can obtain from (11) that $0=J(t)-r(t)-\alpha J(t+1)$, where $J(t)=J(x(t))$.
III-B Critic network
The critic network is a one-hidden-layer multilayer perceptron (MLP) that takes the input as $y(t)=\big(x_{1}(t),\ldots,x_{m}(t)\big)^{T}$ and the output as $\hat{J}(t)$ to approximate $J(t)$. The number of nodes in hidden layer is $N_{h_{c}}$. The weight between input node $i$ and hidden node $j$ is $\hat{w}^{(1)}_{c,ji}(t)$. The weight between hidden node $j$ and output node is $\hat{w}^{(2)}_{c,j}(t)$.
The input to hidden node $j$ is:
$\sigma_{c,j}(t)=\sum_{i=1}^{m}\hat{w}^{(1)}_{c,ji}(t)\,x_{i}(t)$ (12)
The output of hidden node $j$ is
$\phi_{c,j}(t)=\tanh(\sigma_{c,j}(t))$ (13)
where $\tanh(\cdot)$ is the nonlinear activation function.
The output of the critic is
$\hat{J}(t)=\sum_{j=1}^{N_{h_{c}}}\hat{w}^{(2)}_{c,j}(t)\,\phi_{c,j}(t),$ (14)
III-C Actor network
The actor network is a one-hidden-layer multilayer perceptron (MLP) that takes the input as the state $x(t)=\big(x_{1}(t),\ldots,x_{m}(t)\big)^{T}$, and the output as the action. The number of nodes in hidden layer is $N_{h_{a}}$. The weight between input node $i$ and hidden node $j$ is $\hat{w}^{(1)}_{a,ji}(t)$. The weight between hidden node $j$ and output node $k$ is $\hat{w}^{(2)}_{a,kj}(t)$.
The input to hidden node $j$ is
$\sigma_{a,j}(t)=\sum_{i=1}^{m}\hat{w}^{(1)}_{a,ji}(t)\,x_{i}(t).$ (15)
The output of hidden node $j$ is:
$\phi_{a,j}(t)=\tanh(\sigma_{a,j}(t))$ (16)
The hidden to output node is
$\sigma_{a,k}(t)=\sum_{j=1}^{N_{h_{a}}}\hat{w}^{(2)}_{a,kj}(t)\,\phi_{a,j}(t)$ (17)
The output of the output node $k$ is
$u_{k}(t)=\tanh(\sigma_{a,k}(t))$ (18)
III-D Update of the critic network
The critic network is trained to minimize the objective function $E_{c}(t)=\frac{1}{2}e_{c}^{2}(t)$, where $e_{c}(t)=\hat{J}(t)-r(t)-\alpha\hat{J}(t+1)$ is the Bellman error. This definition in ADHDP [Ref] follows a backforward style due to the convenience of policy improvement using state-action value function.
The weights between the input layer and the hidden layer are updated by
$\hat{w}^{(1)}_{c}(t+1)=\hat{w}^{(1)}_{c}(t+1)+\Delta\hat{w}^{(1)}_{c}(t)$ (19)
where $\hat{w}^{(1)}_{c}\in\mathbb{R}^{N_{h_{c}}\times m}$.
According to the gradient descent:
$\Delta\hat{w}^{(1)}_{c,ji}(t)=l_{c}\left[-\frac{\partial E_{c}(t)}{\partial\hat{w}^{(1)}_{c,ji}(t)}\right]$ (20)
which follows from the chain rule
$\frac{\partial E_{c}(t)}{\partial\hat{w}^{(1)}_{c,ji}(t)}$ $=\frac{\partial E_{c}(t)}{\partial e_{c}(t)}\frac{\partial e_{c}(t)}{\partial\hat{J}(t)}\frac{\partial\hat{J}(t)}{\partial\phi_{c,j}(t)}\frac{\partial\phi_{c,j}(t)}{\partial\sigma_{c,j}(t)}\frac{\partial\sigma_{c,j}(t)}{\partial\hat{w}^{(1)}_{c,ji}(t)}$ (21)
$=e_{c}(t)\,\hat{w}^{(2)}_{c,j}(t)\phi_{c,j}^{\prime}(t)y_{i}(t)$
The weights between the hidden layer and the output layer are updated by
$\hat{w}^{(2)}_{c}(t+1)=\hat{w}^{(2)}_{c}(t+1)+\Delta\hat{w}^{(2)}_{c}(t)$ (22)
where $\hat{w}^{(2)}_{c}\in\mathbb{R}^{1\times N_{h_{c}}}$.
According to the gradient descent:
$\Delta\hat{w}^{(2)}_{c,j}(t)=l_{c}\left[-\frac{\partial E_{c}(t)}{\partial\hat{w}^{(2)}_{c,j}(t)}\right]$ (23)
which follows from the chain rule
$\frac{\partial E_{c}(t)}{\partial\hat{w}^{(2)}_{c,j}(t)}=\frac{\partial E_{c}(t)}{\partial e_{c}(t)}\frac{\partial e_{c}(t)}{\partial\hat{J}(t)}\frac{\partial\hat{J}(t)}{\partial\hat{w}^{(2)}_{c,j}(t)}=e_{c}(t)\,\phi_{c,j}(t)$ (24)
III-E Update of the actor network
The actor network is trained according to $\hat{J}(t)$, which is reconstructed from the Bellman equation as $\hat{J}(t)=r(t)+\alpha\hat{J}(t+1)$, rather than a state–action value function $\hat{J}(x(t),u(t))$ as in ADHDP [Ref]. (In this paper, we use an importance-weighted training approach). (Let $U_{c}$ denote the desired ultimate objective function. The quadratic error measure to be minimized is $E_{a}(t)=E_{a1}(t)+\lambda E_{a2}(t)$, $E_{a1}(t)=\frac{1}{2}e_{a}^{2}(t),E_{a2}(t)=\frac{1}{2}e^{2}_{s}(t),e_{a}(t)=r(t)+\alpha\hat{J}(t+1)-U_{c}$, $e_{s}(t)=\frac{1}{2}\|u(t)-u(t+1)\|^{2}$. $e_{s}(t)$ is a temporal smoothness measure weighted by $\lambda\geq 0$. In reinforcement learning, success corresponds to an objective that is zero at each time step. For simplicity of derivation, we assume $U_{c}=0$, i.e. $e_{a}(t)=r(t)+\alpha\hat{J}(t+1)$).
The gradient-descent update rule is $\hat{w}_{a}(t+1)=\hat{w}_{a}(t)+\Delta\hat{w}_{a}(t)$, where $\Delta\hat{w}_{a}(t)=l_{a}\left[-\frac{\partial E_{a}(t)}{\partial\hat{w}_{a}(t)}\right],\ l_{a}>0$.
The weights between the input layer and the hidden layer is updated by
$\hat{w}^{(1)}_{a}(t+1)=\hat{w}^{(1)}_{a}(t+1)+\Delta\hat{w}^{(1)}_{a}(t)$ (25)
where $\hat{w}^{(1)}_{c}\in\mathbb{R}^{N_{h_{c}}\times m}$, and the element-wise increment:
$\Delta\hat{w}^{(1)}_{a,ji}(t)=l_{a}\left[-\frac{\partial E_{a1}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}-\lambda\frac{\partial E_{a2}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}\right],$ (26)
Applying the chain rule:
$\frac{\partial E_{a1}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$ $=\frac{\partial E_{a1}(t)}{\partial e_{a}(t)}\left[\frac{\partial e_{a}(t)}{\partial u(t)}\right]^{T}\frac{\partial u(t)}{\partial\sigma_{a,j}(t)}\frac{\partial\sigma_{a,j}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$ (27)
$=\frac{\partial E_{a1}(t)}{\partial e_{a}(t)}\sum^{n}_{k=1}\frac{\partial e_{a}(t)}{\partial u_{k}(t)}\frac{\partial u_{k}(t)}{\partial\sigma_{a,k}(t)}\frac{\partial\sigma_{a,k}(t)}{\partial\phi_{a,j}(t)}\frac{\partial\phi_{a,j}(t)}{\partial\sigma_{a,j}(t)}\frac{\partial\sigma_{a,j}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$
$=\frac{\partial E_{a1}(t)}{\partial e_{a}(t)}\sum^{n}_{k=1}\left(\frac{\partial r(t)}{\partial u(t)}+\alpha\left[\frac{\partial\hat{J}(t+1)}{\partial x(t+1)}\right]^{T}\frac{\partial x(t+1)}{\partial u(t)}\right)_{k}\frac{\partial u_{k}(t)}{\partial\sigma_{a,k}(t)}\frac{\partial\sigma_{a,k}(t)}{\partial\phi_{a,j}(t)}\frac{\partial\phi_{a,j}(t)}{\partial\sigma_{a,j}(t)}\frac{\partial\sigma_{a,j}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$
$=\frac{\partial E_{a1}(t)}{\partial e_{a}(t)}\sum^{n}_{k=1}\left(2Ru(t)+\alpha\left[\frac{\partial\hat{J}(t+1)}{\partial x(t+1)}\right]^{T}G(t-1)\right)_{k}\frac{\partial u_{k}(t)}{\partial\sigma_{a,k}(t)}\frac{\partial\sigma_{a,k}(t)}{\partial\phi_{a,j}(t)}\frac{\partial\phi_{a,j}(t)}{\partial\sigma_{a,j}(t)}\frac{\partial\sigma_{a,j}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$
$=e_{a}(t)\sum^{n}_{k=1}\left(2Ru(t)+\left[\alpha\hat{w}^{(2)}_{c}(t)\phi^{\prime}(x(t+1))\hat{w}^{(1)}_{c}(t)\right]^{T}G(t-1)\right)_{k}\phi^{\prime}_{a,k}(t)\hat{w}^{(2)}_{a,kj}(t)\phi_{a,j}^{\prime}(x(t))x_{i}(t)$
and
$\frac{\partial E_{a2}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$ $=\frac{\partial E_{a2}(t)}{\partial e_{s}(t)}\left[\frac{\partial e_{s}(t)}{\partial u(t)}\right]^{T}\frac{\partial u(t)}{\partial\sigma_{a,k}(t)}\frac{\partial\sigma_{a,k}(t)}{\partial\phi_{a,j}(t)}\frac{\partial\phi_{a,j}(t)}{\partial\sigma_{a,j}(t)}\frac{\partial\sigma_{a,j}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$ (28)
$=\frac{\partial E_{a2}(t)}{\partial e_{s}(t)}\sum^{n}_{k=1}\frac{\partial e_{s}(t)}{\partial u_{k}(t)}\frac{\partial u_{k}(t)}{\partial\sigma_{a,k}(t)}\frac{\partial\sigma_{a,k}(t)}{\partial\phi_{a,j}(t)}\frac{\partial\phi_{a,j}(t)}{\partial\sigma_{a,j}(t)}\frac{\partial\sigma_{a,j}(t)}{\partial\hat{w}^{(1)}_{a,ji}(t)}$
$=e_{s}(t)\sum^{n}_{k=1}(u_{k}(t)-u_{k}(t+1))\phi^{\prime}_{a,k}(t)\hat{w}^{(2)}_{a,kj}(t)\phi_{a,j}^{\prime}(x(t))x_{i}(t)$
The weight between the hidden layer and the output layer is updated by
$\hat{w}^{(2)}_{a}(t+1)=\hat{w}^{(2)}_{a}(t+1)+\Delta\hat{w}^{(2)}_{a}(t)$ (29)
where $\hat{w}^{(2)}_{a}\in\mathbb{R}^{n\times N_{h_{a}}}$, and the element-wise increment:
$\Delta\hat{w}^{(2)}_{a,kj}(t)=l_{a}\left[-\frac{\partial E^{1}_{a}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}-\lambda\frac{\partial E^{2}_{a}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}\right]$ (30)
Applying the chain rule:
$\frac{\partial E_{a1}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}$ $=\frac{\partial E_{a1}(t)}{\partial e_{a}(t)}\frac{\partial e_{a}(t)}{\partial u_{k}(t)}\frac{\partial u_{k}(t)}{\partial\sigma_{a,k}(t)}\frac{\partial\sigma_{a,k}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}$ (31)
$=\frac{\partial E_{a1}(t)}{\partial e_{a}(t)}\left(\frac{\partial r(t)}{\partial u(t)}+\alpha\left[\frac{\partial\hat{J}(t+1)}{\partial x(t+1)}\right]^{T}\frac{\partial x(t+1)}{\partial u(t)}\right)_{k}\frac{\partial u_{k}(t)}{\partial\sigma_{a,k}(t)}\frac{\partial\sigma_{a,k}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}$
$=e_{a}(t)\left(2Ru(t)+\alpha\left[\frac{\partial\hat{J}(t+1)}{\partial x(t+1)}\right]^{T}G(t-1)\right)_{k}\phi^{\prime}_{a,k}(t)\phi_{a,j}(t)$
$=e_{a}(t)\left(2Ru(t)+\left[\alpha\hat{w}^{(2)}_{c}(t)\phi^{\prime}(x(t+1))\hat{w}^{(1)}_{c}(t)\right]^{T}G(t-1)\right)_{k}\phi^{\prime}_{a,k}(t)\phi_{a,j}(t)$
and
$\frac{\partial E_{a2}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}$ $=\frac{\partial E_{a2}(t)}{\partial e_{s}(t)}\frac{\partial e_{s}(t)}{\partial u_{k}(t)}\frac{\partial u_{k}(t)}{\partial\sigma_{a,k}(t)}\frac{\partial\sigma_{a,k}(t)}{\partial\hat{w}^{(2)}_{a,kj}(t)}$ (32)
$=e_{s}(t)(u_{k}(t)-u_{k}(t+1))\phi^{\prime}_{a,k}(t)\phi_{a,j}(t)$
Remark 1.
IHDP abstracts the control effectiveness matrix $G_{t-1}$ from a state-action function in ADHDP, improving the learning accuracy and efficiency. The matrix $G_{t-1}$ is identified using RLS algorithm online.
IV Convergence Analysis
IV-A Basics of uniformly ultimately bounded stability
Denote the sub-optimal weights by ${w}_{c}^{*},w_{a}^{*}$, defined as $w_{c}^{*}=\arg\min_{\hat{w}_{c}}\left\|\hat{J}(t)-r(t)-\alpha\hat{J}(t+1)\right\|$, and $w_{a}^{*}=\arg\min_{\hat{w}_{a}}\left\|r(t)+\alpha\hat{J}(t+1)\right\|$ if assuming $U_{c}=0$. The estimation errors of the critic and actor weights are defined by $\tilde{w}(t):=\hat{w}(t)-w^{*}$. Then Equations 21, 24, 27,32 define a dynamical system of estimation errors for a nonlinear function $F$:
$\tilde{w}(t+1)=\tilde{w}(t)-F\!\left(\hat{w}(t),\phi(t),\phi(t+1)\right)$ (33)
Definition 1.
A dynamical system is said to be uniformly ultimately bounded (UUB) with ultimate bound $b>0$, if for any $a>0$ and $t_{0}>0$, there exists a positive number $N=N(a,b)$ independent of $t_{0}$, such that $\|\tilde{w}(t)\|\leq b,\text{for all }t\geq N+t_{0}$, whenever $\|\tilde{w}(t_{0})\|\leq a$.
Theorem 1 (UUB of a discrete dynamical system).
If, for system (33), there exists a function $L(\tilde{w}(t),t)$ such that for all $\tilde{w}(t_{0})$ in a compact set $K$, $L(\tilde{w}(t),t)$ is positive definite and the first difference satisfies $\Delta L(\tilde{w}(t),t)<0,\text{for}\;\|\tilde{w}(t_{0})\|>b$, for some $b>0$ such that the $b$-neighborhood of $\tilde{w}(t)$ is contained in $K$, then the system is UUB and the norm of the state is bounded within a neighborhood of $b$.
IV-B Network weights increments
The convergence of critic and actor weights is analyzed using gradient descent algorithm. The requirements on learning rate to stabilize the convergence are analyzed. The ultimate bound is also provided.
Assumption 1.
The suboptimal weights are bounded as $\|w_{c}^{\ast}\|\leq w_{c}^{\max}$, $\|w_{a}^{\ast}\|\leq w_{a}^{\max}$.
Lemma 1.
Under Assumption 1, define $L_{1}(t)=\frac{1}{l_{c}}\operatorname{tr}\!\left[(\tilde{w}_{c}^{(2)}(t))^{T}\tilde{w}_{c}^{(2)}(t)\right]$, and its first difference $\Delta L_{1}(t)=L_{1}(t+1)-L_{1}(t)$ is
$\Delta L_{1}(t)$ $=l_{c}e_{c}^{2}(t)\|\phi_{c}(t)\|^{2}+\|e_{c}(t)-\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-\|\zeta_{c}(t)\|^{2}-e^{2}_{c}(t)$ (34)
where $\zeta_{c}(t)=\tilde{w}_{c}^{(2)}(t)\phi_{c}(t)$ is the approximation error of the critic output.
Proof.
According to (22), (23), (24) and the fact that $w_{c}^{\ast(2)}$ is time-invariant:
$\tilde{w}_{c}^{(2)}(t+1)$ $=\hat{w}_{c}^{(2)}(t+1)-w_{c}^{\ast(2)}=\tilde{w}_{c}^{(2)}(t)-l_{c}e_{c}(t)\phi^{T}_{c}(t)$ (35)
where $\phi_{c}(t)\in\mathbb{R}^{N_{h_{c}}\times 1}$, $e_{c}(t)$ is a scaler. The quadratic weight error is
$\text{tr}\left[\left(\tilde{w}_{c}^{(2)}(t+1)\right)^{T}\tilde{w}_{c}^{(2)}(t+1)\right]=$ $\text{tr}\left[\left(\tilde{w}_{c}^{(2)}(t)-l_{c}e_{c}(t)\phi^{T}_{c}(t)\right)^{T}\left(\tilde{w}_{c}^{(2)}(t)-l_{c}e_{c}(t)\phi^{T}_{c}(t)\right)\right]$ (36)
$=$ $\text{tr}\left[\left(\hat{w}_{c}^{(2)}(t)\right)^{T}\tilde{w}_{c}^{(2)}(t)\right]+\text{tr}\left[l^{2}_{c}e^{2}_{c}(t)\phi_{c}(t)\phi^{T}_{c}(t)\right]$
$\quad-l_{c}e_{c}(t)\text{tr}\left[\left(\tilde{w}_{c}^{(2)}(t)\right)^{T}\phi^{T}_{c}(t)\right]-l_{c}e_{c}(t)\text{tr}\left[\phi_{c}(t)\tilde{w}_{c}^{(2)}(t)\right]$
$=$ $\text{tr}\left[\left(\hat{w}_{c}^{(2)}(t)\right)^{T}\tilde{w}_{c}^{(2)}(t)\right]+\text{tr}\left[l^{2}_{c}e^{2}_{c}(t)\phi_{c}(t)\phi^{T}_{c}(t)\right]$
$\quad-2l_{c}e_{c}(t)\text{tr}\left[\phi_{c}(t)\tilde{w}_{c}^{(2)}(t)\right]$
$=$ $\text{tr}\left[\left(\hat{w}_{c}^{(2)}(t)\right)^{T}\tilde{w}_{c}^{(2)}(t)\right]+l^{2}_{c}e^{2}_{c}(t)\|\phi_{c}(t)\|^{2}$
$\quad-2l_{c}e_{c}(t)\tilde{w}_{c}^{(2)}(t)\phi_{c}(t)$
Since $\tilde{w}_{c}^{(2)}(t)\phi_{c}(t)$ is a scalar, the cross term becomes
$-2l_{c}e_{c}(t)\tilde{w}_{c}^{(2)}(t)\phi_{c}(t)=l_{c}\Bigl(\|e_{c}(t)-\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-\|\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-e_{c}^{2}(t)\Bigr)$ (37)
The difference $\Delta L_{1}(t)$ yields
$\Delta L_{1}(t)$ $=\frac{1}{l_{c}}\Bigl(\operatorname{tr}[(\tilde{w}_{c}^{(2)}(t+1))^{T}\tilde{w}_{c}^{(2)}(t+1)]-\operatorname{tr}[(\tilde{w}_{c}^{(2)}(t))^{T}\tilde{w}_{c}^{(2)}(t)]\Bigr)$ (38)
$=l_{c}e_{c}^{2}(t)\|\phi_{c}(t)\|^{2}+\|e_{c}(t)-\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-\|\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-e^{2}_{c}(t)$
$=\underbrace{l_{c}e_{c}^{2}(t)\|\phi_{c}(t)\|^{2}+\|e_{c}(t)-\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-\|\zeta_{c}(t)\|^{2}-e^{2}_{c}(t)}_{\Delta L_{1}(t)}$
where $\zeta_{c}(t)=\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)$. ∎
Lemma 2.
Under Assumption 1, define $L_{2}(t)=\frac{1}{l_{a}\gamma_{1}}\operatorname{tr}\!\left[(\tilde{w}_{a}^{(2)}(t))^{T}\tilde{w}_{a}^{(2)}(t)\right]$, and its first difference $\Delta L_{2}(t)=L_{1}(t+1)-L_{1}(t)$ is bounded by
$\Delta L_{2}(t)=\frac{1}{l_{a}\gamma_{1}}\mathrm{tr}\!\left(\tilde{w}^{(2)}_{a}(t+1)^{T}\tilde{w}^{(2)}_{a}(t+1)-\tilde{w}^{(2)}_{a}(t)^{T}\tilde{w}^{(2)}_{a}(t)\right)$ (39)
$\leq\frac{1}{\gamma_{1}}\left(2l_{a}e^{2}_{a}(t)\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}+4\|r(t)\|^{2}+8\alpha^{2}\left\|\zeta_{c}(t+1)\right\|^{2}+8\alpha^{2}\left\|w^{*(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\|C(t)\|^{2}\|\zeta_{a}(t)\|^{2}-e^{2}_{a}(t)\right)$
$+\frac{\lambda}{\gamma_{1}}\left(2\lambda l_{a}e^{2}_{s}(t)\|U(t)\|^{2}\|\phi_{a}(t)\|^{2}+2e^{2}_{s}(t)+\left\|U^{T}(t)\zeta_{a}(t)\right\|^{2}-e^{2}_{s}(t)\right)$
where $\zeta_{a}(t)=\tilde{w}_{a}^{(2)}(t)\phi_{a}(t)$.
Proof.
According to (29), (30), (32) and the fact that $w^{*(2)}_{a}$ is time-invariant:
$\tilde{w}^{(2)}_{a}(t+1)$ $=\hat{w}^{(2)}_{a}(t+1)-w^{*(2)}_{a}=\tilde{w}^{(2)}_{a}(t)-l_{a}e_{a}(t)C(t)\phi^{T}_{a}(t)-\lambda l_{a}e_{s}(t)U(t)\phi^{T}_{a}(t)$ (40)
where $\phi_{a}\in\mathbb{R}^{N_{a_{h}}\times 1}$, $C(t)\in\mathbb{R}^{n\times 1}$ and its element $C_{k}(t)=\left(2Ru(t)+\left[\alpha\hat{w}^{(2)}_{c}(t)\phi^{\prime}(x(t+1))\hat{w}^{(1)}_{c}(t)\right]^{T}G(t-1)\right)_{k}\phi^{\prime}_{a,k}(t)$. $U(t)\in\mathbb{R}^{n\times 1}$ and its element $U_{k}(t)=\left(u_{k}(t)-u_{k}(t+1)\right)\phi^{\prime}_{a,k}(t)$.
The trace of the quadratic weight error is
$\text{tr}\left[\left(\tilde{w}^{(2)}_{a}(t+1)\right)^{T}\tilde{w}^{(2)}_{a}(t+1)\right]$ (41)
$=$ $\text{tr}\left[\left(\tilde{w}^{(2)}_{a}(t)-l_{a}e_{a}(t)C(t)\phi^{T}_{a}(t)-\lambda l_{a}e_{s}(t)U(t)\phi^{T}_{a}(t)\right)^{T}\left(\tilde{w}^{(2)}_{a}(t)-l_{a}e_{a}(t)C(t)\phi^{T}_{a}(t)-\lambda l_{a}e_{s}(t)U(t)\phi^{T}_{a}(t)\right)\right]$
$=$ $\text{tr}\left[\left(\tilde{w}^{(2)}_{a}(t)\right)^{T}\tilde{w}^{(2)}_{a}(t)\right]+\text{tr}\left[\left(l_{a}e_{a}(t)C(t)\phi^{T}_{a}(t)+\lambda l_{a}e_{s}(t)U(t)\phi^{T}_{a}(t)\right)^{T}\left(l_{a}e_{a}(t)C(t)\phi^{T}_{a}(t)+\lambda l_{a}e_{s}(t)U(t)\phi^{T}_{a}(t)\right)\right]$
$\qquad-l_{a}e_{a}(t)\text{tr}\left[\left(\tilde{w}^{(2)}_{a}(t)\right)^{T}C(t)\phi^{T}_{a}(t)\right]-l_{a}e_{a}(t)\text{tr}\left[\phi_{a}(t)C^{T}(t)\tilde{w}^{(2)}_{a}(t)\right]$
$\qquad-\lambda l_{a}e_{s}(t)\text{tr}\left[\left(\tilde{w}^{(2)}_{a}(t)\right)^{T}U(t)\phi^{T}_{a}(t)\right]-\lambda l_{a}e_{s}(t)\text{tr}\left[\phi_{a}(t)U^{T}(t)\tilde{w}^{(2)}_{a}(t)\right]$
$=$ $\text{tr}\left[\left(\tilde{w}^{(2)}_{a}(t)\right)^{T}\tilde{w}^{(2)}_{a}(t)\right]+\text{tr}\left[l^{2}_{a}e^{2}_{a}(t)\phi_{a}(t)C^{T}(t)C(t)\phi^{T}_{a}(t)\right]+\text{tr}\left[\lambda^{2}l^{2}_{a}e^{2}_{s}(t)\phi_{a}(t)U^{T}(t)U(t)\phi^{T}_{a}(t)\right]$
$\quad+2\lambda l^{2}_{a}e_{a}(t)e_{s}(t)\text{tr}\left[\phi_{a}(t)C^{T}(t)U(t)\phi_{a}^{T}(t)\right]-2l_{a}e_{a}(t)\text{tr}\left[\phi_{a}(t)C^{T}(t)\tilde{w}^{(2)}_{a}(t)\right]-2\lambda l_{a}e_{s}(t)\text{tr}\left[\phi_{a}(t)U^{T}(t)\tilde{w}^{(2)}_{a}(t)\right]$
$=$ $\text{tr}\left[\left(\tilde{w}^{(2)}_{a}(t)\right)^{T}\tilde{w}^{(2)}_{a}(t)\right]+l^{2}_{a}e^{2}_{a}(t)\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}+\lambda^{2}l^{2}_{a}e^{2}_{s}(t)\|U(t)\|^{2}\|\phi_{a}(t)\|^{2}$
$\quad+\underbrace{2\lambda l^{2}_{a}e_{a}(t)e_{s}(t)\text{tr}\left[\phi_{a}(t)C^{T}(t)U(t)\phi_{a}^{T}(t)\right]}_{\text{term1}}-2l_{a}e_{a}(t)\text{tr}\left[\phi_{a}(t)C^{T}(t)\tilde{w}^{(2)}_{a}(t)\right]-\underbrace{2\lambda l_{a}e_{s}(t)\text{tr}\left[\phi_{a}(t)U^{T}(t)\tilde{w}^{(2)}_{a}(t)\right]}_{\text{term 2}}$
Use Cauchy-Schwarz inequality:
$\text{tr}\left[\phi_{a}(t)C^{T}(t)U(t)\phi_{a}^{T}(t)\right]$ $=\text{tr}\left[C^{T}(t)U(t)\left(\phi^{T}_{a}(t)\phi_{a}(t)\right)\right]$ (42)
$=\text{tr}\left[C^{T}(t)U(t)\right]\text{tr}[\phi^{T}(t)\phi(t)]$
$\leq\|C(t)\|\|U(t)\|\|\phi(t)\|^{2}$
Therefore,
$2\lambda l^{2}_{a}e_{a}(t)e_{s}(t)\text{tr}\left[\phi_{a}(t)C^{T}(t)U(t)\phi_{a}^{T}(t)\right]$ $\leq 2\lambda l^{2}_{a}e_{a}(t)e_{s}(t)\|C(t)\|\|U(t)\|\|\phi(t)\|^{2}$ (43)
$\leq l^{2}_{a}e^{2}_{a}(t)\|C(t)\|^{2}\|\phi(t)\|^{2}+\lambda^{2}l^{2}_{a}e^{2}_{s}(t)\|U(t)\|^{2}\|\phi(t)\|^{2}$
Consider the cross term, and use $\text{tr}(ABC)=\text{tr}(CAB)$:
$-2l_{a}e_{a}(t)\text{tr}\left(\phi_{a}(t)C^{T}(t)\tilde{w}^{(2)}_{a}(t)\right)$ $=-2l_{a}e_{a}(t)\text{tr}\left(C^{T}(t)\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)\right)$ (44)
$=-2l_{a}e_{a}(t)C^{T}(t)\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)$
$\leq l_{a}\left(\left\|e_{a}(t)-C^{T}(t)\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)\right\|^{2}-\left\|C^{T}(t)\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)\right\|^{2}-e^{2}_{a}(t)\right)$
$\leq l_{a}\left(\left\|e_{a}(t)-C^{T}(t)\zeta_{a}(t)\right\|^{2}-\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}-e^{2}_{a}(t)\right)$
where $\zeta_{a}(t)=\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)$.
Similarly,
$-2\lambda l_{a}e_{s}(t)\text{tr}\left(\phi_{a}(t)U^{T}(t)\tilde{w}^{(2)}_{a}(t)\right)$ $=-2\lambda l_{a}e_{s}(t)\text{tr}\left(U^{T}(t)\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)\right)$ (45)
$=-2\lambda l_{a}e_{s}(t)U^{T}(t)\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)$
$\leq\lambda l_{a}\left(2e^{2}_{s}(t)+\|U^{T}(t)\tilde{w}^{(2)}_{a}(t)\phi_{a}(t)\|^{2}-e^{2}_{s}(t)\right)$
Notice that $(a-b)^{2}-b^{2}\leq 2a^{2}+b^{2}$, and use $\|C^{T}(t)\zeta_{a}(t)\|^{2}\leq\|C^{T}(t)\|^{2}\|\zeta_{a}(t)\|^{2}$ by Cauchy–Schwarz inequality:
$\left\|e_{a}(t)-C^{T}(t)\zeta_{a}(t)\right\|^{2}-\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$ (46)
$\quad\leq 2\left\|e_{a}(t)\right\|^{2}+\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$
$\quad\leq 2\left\|r(t)+\alpha\hat{J}(t+1)\right\|^{2}+\|C^{T}(t)\zeta_{a}(t)\|^{2}$
$\quad\leq 2\left(\left\|r(t)\right\|+\alpha\left\|\hat{J}(t+1)\right\|\right)^{2}+\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$
$\quad\leq 4\|r(t)\|^{2}+4\alpha^{2}\left\|\hat{w}^{(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$
$\quad\leq 4\|r(t)\|^{2}+4\alpha^{2}\left\|(\tilde{w}^{(2)}_{c}+w^{*(2)}_{c})\phi_{c}(t+1)\right\|^{2}+\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$
$\quad\leq 4\|r(t)\|^{2}+8\alpha^{2}\left\|\tilde{w}^{(2)}_{c}\phi_{c}(t+1)\right\|^{2}+8\alpha^{2}\left\|w^{*(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$
$\quad\leq 4\|r(t)\|^{2}+8\alpha^{2}\left\|\zeta_{c}(t+1)\right\|^{2}+8\alpha^{2}\left\|w^{*(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\left\|C^{T}(t)\zeta_{a}(t)\right\|^{2}$
$\quad\leq 4\|r(t)\|^{2}+8\alpha^{2}\left\|\zeta_{c}(t+1)\right\|^{2}+8\alpha^{2}\left\|w^{*(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\left\|C(t)\|^{2}\|\zeta_{a}(t)\right\|^{2}$
where $\zeta_{c}(t+1)=\tilde{w}^{(2)}_{c}\phi_{c}(t+1)$.
Therefore, the first difference is
$\Delta L_{2}(t)=\frac{1}{l_{a}\gamma_{1}}\mathrm{tr}\!\left(\tilde{w}^{(2)}_{a}(t+1)^{T}\tilde{w}^{(2)}_{a}(t+1)-\tilde{w}^{(2)}_{a}(t)^{T}\tilde{w}^{(2)}_{a}(t)\right)$ (47)
$\leq\frac{1}{\gamma_{1}}\left(2l_{a}e^{2}_{a}(t)\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}+4\|r(t)\|^{2}+8\alpha^{2}\left\|\zeta_{c}(t+1)\right\|^{2}+8\alpha^{2}\left\|w^{*(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\|C(t)\|^{2}\|\zeta_{a}(t)\|^{2}-e^{2}_{a}(t)\right)$
$+\frac{\lambda}{\gamma_{1}}\left(2\lambda l_{a}e^{2}_{s}(t)\|U(t)\|^{2}\|\phi_{a}(t)\|^{2}+2e^{2}_{s}(t)+\left\|U^{T}(t)\zeta_{a}(t)\right\|^{2}-e^{2}_{s}(t)\right)$
Remark 2.
Note that $\Delta L_{2}$ depends on $C(t)$, which incorporates the control-effectiveness matrix $G_{t-1}$. Thus, a system with high control effectiveness, i.e. a more “sensitive” system, may exhibit a larger bound. Intuitively, this matches our expectation.
Remark 3.
$C(t)$ is defined differently from [Ref] given the objective function to improve the policy, i.e., $r(t)+\hat{J}(t+1)$ in IHDP and $\hat{J}(t)$ in ADHDP.
Remark 4.
If we introduce the normalization $\|(\hat{w}_{c}^{(2)}(t))^{T}C(t)\|^{2}=1$ and fix the input-layer weights, then Lemmas 1–2 yield the results of Liu et al. (2012).
Lemma 3.
Under Assumption 1, define $L_{3}(t)=\frac{1}{l_{c}\gamma_{2}}\text{tr}\left[(\tilde{w}_{c}^{(1)}(t))^{T}\,\tilde{w}_{c}^{(1)}(t)\right]$, and its difference $\Delta L_{3}(t)=L_{3}(t+1)-L_{3}(t)$ is bounded by
$\Delta L_{3}(t)\leq\frac{1}{\gamma_{2}}\Big(l_{c}e^{2}_{c}(t)\|a(t)\|^{2}\|y(t)\|^{2}+\|a(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+e^{2}_{c}(t)\Big)$ (48)
where $\gamma_{2}>0$ is a weighting factor and $a(t)\in\mathbb{R}^{N_{h_{c}}\times 1}$ is a vector with $a_{j}(t)=\hat{w}^{(2)}_{c,j}(t)\phi^{\prime}_{j}(t)$.
Proof.
Consider the update rule (21), the weight error is
$\tilde{w}^{(1)}_{c}(t+1)=\hat{w}^{(1)}_{c}(t+1)-w_{c}^{*(1)}=\tilde{w}^{(1)}_{c}(t)-l_{c}e_{c}(t)B(t)$ (49)
where $B(t)\in\mathbb{R}^{N_{h_{c}}\times m}$ and its element $B_{ji}(t)=\hat{w}^{(2)}_{c,j}(t)\phi_{j}^{\prime}(t)y_{i}(t)$, then $B(t)=a(t)y^{T}(t)$.
The quadratic weight error is
$\text{tr}\left[\tilde{w}^{(1)}_{c}(t+1)^{T}\tilde{w}^{(1)}_{c}(t+1)\right]$ $=\text{tr}\left[\left(\tilde{w}^{(1)}_{c}(t)-l_{c}e_{c}(t)B(t)\right)^{T}\left(\tilde{w}^{(1)}_{c}(t)-l_{c}e_{c}(t)B(t)\right)\right]$ (50)
$=\text{tr}\left[(\tilde{w}^{(1)}_{c}(t))^{T}\tilde{w}^{(1)}_{c}(t)\right]+\text{tr}\left[l_{c}^{2}e^{2}_{c}(t)B^{T}(t)B(t)\right]$
$\qquad-l_{c}e_{c}(t)\left[\left(\tilde{w}^{(1)}_{c}(t)\right)^{T}B(t)\right]-l_{c}e_{c}(t)\text{tr}\left[B^{T}(t)\tilde{w}^{(1)}_{c}(t)\right]$
$=\text{tr}\left[(\tilde{w}^{(1)}_{c}(t))^{T}\tilde{w}^{(1)}_{c}(t)\right]+\text{tr}\left[l_{c}^{2}e^{2}_{c}(t)B^{T}(t)B(t)\right]$
$\qquad-2l_{c}e_{c}(t)\text{tr}\left[B^{T}(t)\tilde{w}^{(1)}_{c}(t)\right]$
$=\text{tr}\left[(\tilde{w}^{(1)}_{c}(t))^{T}\tilde{w}^{(1)}_{c}(t)\right]+\text{tr}\left[l_{c}^{2}e^{2}_{c}(t)y(t)a^{T}(t)a(t)y^{T}(t)\right]$
$\qquad-2l_{c}e_{c}(t)\text{tr}\left[B^{T}(t)\tilde{w}^{(1)}_{c}(t)\right]$
$=\text{tr}\left[(\tilde{w}^{(1)}_{c}(t))^{T}\tilde{w}^{(1)}_{c}(t)\right]+l_{c}^{2}e^{2}_{c}(t)\|a(t)\|^{2}\|y(t)\|^{2}$
$\qquad-2l_{c}e_{c}(t)\text{tr}\left[B^{T}(t)\tilde{w}^{(1)}_{c}(t)\right]$
where by $\text{tr}(ABC)=\text{tr}(BCA)$:
$\text{tr}\left[B^{T}(t)\tilde{w}^{(1)}_{c}(t)\right]=$ $\text{tr}\left[y(t)a^{T}(t)\tilde{w}^{(1)}_{c}(t)\right]$ (51)
$=$ $\text{tr}\left[a^{T}(t)\tilde{w}^{(1)}_{c}(t)y(t)\right]$
$=$ $a^{T}(t)\tilde{w}^{(1)}_{c}(t)y(t)$
Use inequality $2ab\leq a^{2}+b^{2}$, the cross term becomes
$-2l_{c}e_{c}(t)a^{T}(t)\tilde{w}^{(1)}_{c}(t)y(t)$ $\leq l_{c}\left(\|a^{T}(t)\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+e^{2}_{c}(t)\right)$ (52)
$\leq l_{c}\left(\|a^{T}(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+e^{2}_{c}(t)\right)$
$\leq l_{c}\left(\|a(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+e^{2}_{c}(t)\right)$
Therefore, the first difference can be bounded by
$\Delta L_{3}(t)\leq\underbrace{\frac{1}{\gamma_{2}}\Big(l_{c}e^{2}_{c}(t)\|a(t)\|^{2}\|y(t)\|^{2}+\|a(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+e^{2}_{c}(t)\Big)}_{\Delta L_{3}(t)}$ (53)
Lemma 4.
Under Assumption 1, define $L_{4}(t)=\frac{1}{l_{a}\gamma_{3}}\operatorname{tr}\!\left[\tilde{w}^{(1)}_{a}(t)^{T}\tilde{w}^{(1)}_{a}(t)\right]$, and its first difference $\Delta L_{3}(t)=L_{3}(t+1)-L_{3}(t)$ is bounded by
$\Delta L_{4}(t)\leq\frac{1}{\gamma_{3}}\Big(l_{a}e^{2}_{a}(t)\left\|C^{T}(t)D(t)\right\|^{2}\|x(t)\|^{2}+e^{2}_{a}(t)+\left\|C^{T}(t)D(t)\right\|^{2}\left\|\tilde{w}^{(1)}_{a}(t)\,x(t)\right\|^{2}\Big)$ (54)
where $\gamma_{3}>0$ is a weighting factor, and $D(t)\in\mathbb{R}^{n\times N_{h_{a}}}$, and its element $D_{kj}(t)=\hat{w}^{(2)}_{a,kj}(t)\sigma^{\prime}_{j}(x(t))$.
Proof.
According to (27), the weight errors are
$\tilde{w}^{(1)}_{a}(t+1)$ $=\hat{w}^{(1)}_{a}(t+1)-w_{a}^{*(1)}$ (55)
$=\tilde{w}^{(1)}_{a}(t)-l_{a}e_{a}(t)D^{T}(t)C(t)x^{T}(t)-\lambda l_{a}e_{s}(t)D^{T}(t)U(t)x^{T}$
and
$\text{tr}\left[\left(\tilde{w}^{(1)}_{a}(t+1)\right)^{T}\tilde{w}^{(1)}_{a}(t+1)\right]$ (56)
$=$ $\text{tr}\Big[\left(\tilde{w}^{(1)}_{a}(t)-l_{a}e_{a}(t)D^{T}(t)C(t)x^{T}(t)-\lambda l_{a}e_{s}(t)D^{T}(t)U(t)x^{T}(t)\right)^{T}$
$\qquad\left(\tilde{w}^{(1)}_{a}(t)-l_{a}e_{a}(t)D^{T}(t)C(t)x^{T}(t)-\lambda l_{a}e_{s}(t)D^{T}(t)U(t)x^{T}(t)\right)\Big]$
$=$ $\text{tr}\left[\left(\tilde{w}^{(1)}_{a}(t)\right)^{T}\tilde{w}^{(1)}_{a}(t)\right]+\text{tr}\Big[\left(l_{a}e_{a}(t)D^{T}(t)C(t)x^{T}(t)+\lambda l_{a}e_{s}(t)D^{T}(t)U(t)x^{T}(t)\right)^{T}$
$\qquad\qquad\qquad\qquad\qquad\qquad\left(l_{a}e_{a}(t)D^{T}(t)C(t)x^{T}(t)+\lambda l_{a}e_{s}(t)D^{T}(t)U(t)x^{T}(t)\right)\Big]$
$\qquad-l_{a}(t)\text{tr}\left[\left(\tilde{w}^{(1)}_{a}(t)\right)^{T}\left(e_{a}(t)D^{T}(t)C(t)x^{T}(t)+\lambda e_{s}(t)D^{T}(t)U(t)x^{T}(t)\right)\right]$
$\qquad-l_{a}\text{tr}\left[\left(e_{a}(t)x(t)C^{T}(t)D(t)+\lambda e_{s}(t)x(t)U^{T}(t)D(t)\right)\tilde{w}^{(1)}_{a}(t)\right]$
$=$ $\text{tr}\left[\left(\tilde{w}^{(1)}_{a}(t)\right)^{T}\tilde{w}^{(1)}_{a}(t)\right]+l^{2}_{a}e^{2}_{a}(t)\text{tr}\left[x(t)C^{T}(t)D(t)D^{T}(t)C(t)x^{T}(t)\right]+\lambda^{2}l^{2}_{a}e^{2}_{s}(t)\text{tr}\left[x(t)U^{T}(t)D(t)D^{T}(t)U(t)x^{T}(t)\right]$
$\qquad+2\lambda l^{2}_{a}e_{s}(t)e_{a}(t)\text{tr}\left[x(t)U^{T}(t)D(t)D^{T}(t)C(t)x^{T}(t)\right]-2l_{a}e_{a}(t)\text{tr}\left[x(t)C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)\right]-2\lambda l_{a}e_{s}(t)\text{tr}\left[x(t)U^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)\right]$
$=$ $\text{tr}\left[\left(\tilde{w}^{(1)}_{a}(t)\right)^{T}\tilde{w}^{(1)}_{a}(t)\right]+l^{2}_{a}e^{2}_{a}(t)\|C^{T}(t)D(t)\|^{2}\|x(t)\|^{2}+\lambda^{2}l^{2}_{a}e^{2}_{s}(t)\|U^{T}(t)D(t)\|^{2}\|x(t)\|^{2}$
$\qquad+2\lambda l^{2}_{a}e_{s}(t)e_{a}(t)\text{tr}\left[x(t)U^{T}(t)D(t)D^{T}(t)C(t)x^{T}(t)\right]-2l_{a}e_{a}(t)\text{tr}\left[x(t)C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)\right]-2\lambda l_{a}e_{s}(t)\text{tr}\left[x(t)U^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)\right]$
Using the cyclic property of the trace $\text{tr}(ABC)=\text{tr}(BCA)$ and the fact that $C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)$ and $C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)$ are scalars:
$\text{tr}\left[x(t)C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)\right]$ $=\text{tr}\left[C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)\right]$ (57)
$=C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)$
$\text{tr}\left[x(t)U^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)\right]$ $=\text{tr}\left[U^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)\right]$ (58)
$=U^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)$
The cross term can be transformed into the form using $2ab\leq a^{2}+b^{2}$ and Cauchy–Schwarz inequality:
$-2l_{a}e_{a}(t)C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)$ $\leq l_{a}\left(e^{2}_{a}(t)+\left\|C^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)\right\|^{2}\right)$ (59)
$\leq l_{a}\left(e^{2}_{a}(t)+\|C^{T}(t)D(t)\|^{2}\|\tilde{w}^{(1)}_{a}(t)x(t)\|^{2}\right)$
$-2l_{a}e_{s}(t)U^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)$ $\leq l_{a}\left(e^{2}_{s}(t)+\left\|U^{T}(t)D(t)\tilde{w}^{(1)}_{a}(t)x(t)\right\|^{2}\right)$ (60)
$\leq l_{a}\left(e^{2}_{s}(t)+\|U^{T}(t)D(t)\|^{2}\|\tilde{w}^{(1)}_{a}(t)x(t)\|^{2}\right)$
And
$\text{tr}\left[x(t)U^{T}(t)D(t)D^{T}(t)C(t)x^{T}(t)\right]$ $=\text{tr}\left[U^{T}(t)D(t)D^{T}(t)C(t)x^{T}(t)x(t)\right]$ (61)
$=\text{tr}\left[U^{T}(t)D(t)D^{T}(t)C(t)\right]\text{tr}\left[x^{T}(t)x(t)\right]$
$=\text{tr}\left[U^{T}(t)D(t)D^{T}(t)C(t)\right]\|x(t)\|^{2}$
$\leq\|U^{T}(t)D(t)\|\|D^{T}(t)C(t)\|\|x(t)\|^{2}$
$\leq\|U^{T}(t)D(t)\|\|C^{T}(t)D(t)\|\|x(t)\|^{2}$
which yields
$2\lambda l^{2}_{a}e_{s}(t)e_{a}(t)\text{tr}\left[x(t)U^{T}(t)D(t)D^{T}(t)C(t)x^{T}(t)\right]$ $\leq 2\lambda l^{2}_{a}e_{s}(t)e_{a}(t)\|U^{T}(t)D(t)\|\|C^{T}(t)D(t)\|\|x(t)\|^{2}$ (62)
$\leq 2\lambda l^{2}_{a}e_{s}(t)e_{a}(t)\|U^{T}(t)D(t)\|\|C^{T}(t)D(t)\|\|x(t)\|^{2}$
$\leq l^{2}_{a}\left(e^{2}_{a}(t)\|C^{T}(t)D(t)\|^{2}\|x(t)\|^{2}+\lambda^{2}e^{2}_{s}\|U^{T}(t)D(t)\|^{2}\|x(t)\|^{2}\right)$
Finally, we can obtain the upper bound
$\Delta L_{4}(t)$ $\leq\frac{1}{\gamma_{3}}\Big(2l_{a}e^{2}_{a}(t)\left\|C^{T}(t)D(t)\right\|^{2}\|x(t)\|^{2}+e^{2}_{a}(t)+\left\|C^{T}(t)D(t)\right\|^{2}\left\|\tilde{w}^{(1)}_{a}(t)\,x(t)\right\|^{2}\Big)$ (63)
$+\frac{\lambda}{\gamma_{3}}\Big(2\lambda l_{a}e^{2}_{s}(t)\left\|U^{T}(t)D(t)\right\|^{2}\|x(t)\|^{2}+e^{2}_{s}(t)+\left\|U^{T}(t)D(t)\right\|^{2}\left\|\tilde{w}^{(1)}_{a}(t)\,x(t)\right\|^{2}\Big)$
IV-C Convergence of the weights errors
We introduce the Lyapunov candidate
$L(t)=L_{1}(t)+L_{2}(t)+L_{3}(t)+L_{4}(t)$ (64)
Theorem 2 (Main Theorem).
Let the weights of the critic and actor be updated by gradient descent algorithm, and assume that the reward signal is a bounded semidefinite function. Under Assumption 1, the estimation errors $w_{a}^{\ast}-\hat{w}_{a}(t)$ and $w_{c}^{\ast}-\hat{w}_{c}(t)$ are uniformly ultimately bounded (UUB) if
$l_{c}<\min_{t}\frac{\gamma_{2}-1}{\gamma_{2}\big(\|\phi_{c}(t)\|^{2}+\frac{1}{\gamma_{2}}\|a(t)\|^{2}\|y(t)\|^{2}\big)}$ (65)
$l_{a}<\min_{t}\frac{\gamma_{3}-\gamma_{1}}{\gamma_{3}\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}+\gamma_{1}\|D(t)C^{T}\|^{2}\|x(t)\|^{2}}$ (66)
Proof of Theorem 2.
Collect all terms of $\Delta L(t)$ based on the results of Lemmas 1–4, $\Delta L(t)$ is bounded by
$\Delta L(t)\leq$ $(\Delta L_{1})=l_{c}\underbrace{e_{c}^{2}(t)}_{2}\|\phi_{c}(t)\|^{2}+\|e_{c}(t)-\tilde{w}^{(2)}_{c}(t)\phi_{c}(t)\|^{2}-\underbrace{\|\zeta_{c}(t)\|^{2}}_{1}-\underbrace{e^{2}_{c}(t)}_{2}$ (67)
$(\Delta L_{2})+\frac{1}{\gamma_{1}}\left(2l_{a}\underbrace{e^{2}_{a}(t)}_{3}\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}+4\|r(t)\|^{2}+8\alpha^{2}\underbrace{\left\|\zeta_{c}(t+1)\right\|^{2}}_{1}+8\alpha^{2}\left\|w^{*(2)}_{c}\phi_{c}(t+1)\right\|^{2}+\|C(t)\|^{2}\|\zeta_{a}(t)\|^{2}-\underbrace{e^{2}_{a}(t)}_{3}\right)$ (68)
$\quad\qquad+\frac{\lambda}{\gamma_{1}}\left(2\lambda l_{a}\underbrace{e^{2}_{s}(t)}_{4}\|U(t)\|^{2}\|\phi_{a}(t)\|^{2}+2e^{2}_{s}+\left\|D^{T}(t)\zeta_{a}(t)\right\|^{2}-\underbrace{e^{2}_{s}(t)}_{4}\right)$ (69)
$(\Delta L_{3})+\frac{1}{\gamma_{2}}\Big(l_{c}\underbrace{e^{2}_{c}(t)}_{2}\|a(t)\|^{2}\|y(t)\|^{2}+\|a(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+\underbrace{e^{2}_{c}(t)}_{2}\Big)$ (70)
$\Delta L_{4}(t)+\frac{1}{\gamma_{3}}\Big(2l_{a}\underbrace{e^{2}_{a}(t)}_{3}\left\|C^{T}(t)D(t)\right\|^{2}\|x(t)\|^{2}+\underbrace{e^{2}_{a}(t)}_{3}+\left\|C^{T}(t)D(t)\right\|^{2}\left\|\tilde{w}^{(1)}_{a}(t)\,x(t)\right\|^{2}\Big)$ (71)
$\qquad\quad+\frac{\lambda}{\gamma_{3}}\Big(2\lambda l_{a}\underbrace{e^{2}_{s}(t)}_{4}\left\|U^{T}(t)D(t)\right\|^{2}\|x(t)\|^{2}+\underbrace{e^{2}_{s}(t)}_{4}+\left\|U^{T}(t)D(t)\right\|^{2}\left\|\tilde{w}^{(1)}_{a}(t)\,x(t)\right\|^{2}\Big)$ (72)
After combining like terms, the expression simplifies to
$\Delta L(t)\leq$ $\underbrace{-\Big(\|\zeta_{c}(t)\|^{2}-\frac{8\alpha^{2}}{\gamma_{1}}\|\zeta_{c}(t+1)\|^{2}\Big)}_{1}-\underbrace{\Big(1-l_{c}\|\phi_{c}(t)\|^{2}-\frac{l_{c}}{\gamma_{2}}\|a(t)\|^{2}\|y(t)\|^{2}-\frac{1}{\gamma_{2}}\Big)e^{2}_{c}(t)}_{2}$ (73)
$-\underbrace{\Big(\frac{1}{\gamma_{1}}-\frac{2l_{a}}{\gamma_{1}}\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}-\frac{2l_{a}}{\gamma_{3}}\|C(t)^{T}D(t)\|^{2}\|x(t)\|^{2}-\frac{1}{\gamma_{3}}\Big)e^{2}_{a}(t)}_{3}$
$-\underbrace{\left(\frac{\lambda}{\gamma_{1}}-\frac{2\lambda^{2}l_{a}}{\gamma_{1}}\|U(t)\|^{2}\|\phi_{a}(t)\|^{2}-\frac{2\lambda^{2}l_{a}}{\gamma_{3}}\|U^{T}(t)D(t)\|^{2}\|x(t)\|^{2}-\frac{\lambda}{\gamma_{3}}\right)e^{2}_{s}(t)}_{4}$
$+\big\|w^{*(2)}_{c}(t)\phi_{c}(t)-r(t)-\alpha\hat{w}^{(2)}_{c}(t)\phi(t+1)\|^{2}+4\|r(t)\|^{2}+\frac{8\alpha^{2}}{\gamma_{1}}\|w_{c}^{*(2)}\phi_{c}(t+1)\|^{2}+\frac{1}{\gamma_{1}}\|C(t)\|^{2}\|\zeta_{a}(t)\|^{2}$
$+2\frac{\lambda}{\gamma_{1}}e^{2}_{s}+\frac{\lambda}{\gamma_{1}}\|D^{T}(t)\zeta_{a}(t)\|^{2}+\frac{\lambda}{\gamma_{3}}\|U^{T}(t)D(t)\|^{2}\|\tilde{w}^{(1)}_{a}x(t)\|^{2}$
$+\frac{1}{\gamma_{2}}\|a(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}$
$+\frac{1}{\gamma_{3}}\|C^{T}(t)D(t)\|^{2}\|\tilde{w}^{(1)}_{a}(t)x(t)\|^{2}$
Remark 5.
The boundedness of $e_{s}(t)$ depends on the time step $T$, which can be small when using fast sampling strategy.
To ensure the negativity of the term 1:
$\frac{\gamma_{1}}{8\alpha^{2}}$ $>\frac{\|\zeta_{c}(t+1)\|^{2}}{\|\zeta_{c}(t)\|^{2}}$ (74)
$=\frac{\|\tilde{w}^{(2)}(t)\|^{2}\Big(\|\phi(x(t))\|^{2}+\|L\Delta x(t)\|^{2}\Big)}{\|\tilde{w}^{(2)}(t)\|^{2}\|\phi(x(t))\|^{2}}$
$=\frac{\phi(x(t))\|^{2}+\|L\Delta x(t)\|^{2}}{\|\phi(x(t))\|^{2}}$
$=1+\frac{\|L\Delta x(t)\|^{2}}{\|\phi(x(t))\|^{2}}$
where $\Delta x(t)=x(t+1)-x(t)$. The term $\frac{\|L\Delta x(t)\|^{2}}{\|\phi(x(t))\|^{2}}$ is the effect of the state transition from $x(t),x(t+1)$ on input to hidden layer of the critic, which can be proved to be upper bounded if $x(t),w^{(1)}_{c}(t)$ are upper bounded:
$\frac{\|L\Delta x(t)\|^{2}}{\|\phi(x(t))\|^{2}}\leq$ $\frac{\|L\|^{2}\|\Delta x(t)\|^{2}}{\|\tanh(w^{(1)}_{c}(t)x(t))\|^{2}}$ (75)
$\leq$ $\frac{\|w^{(1)}_{c}(t)x(t)\|^{2}{}\|\Delta x(t)\|^{2}}{\|\tanh(w^{(1)}_{c}(t)x(t))\|^{2}}$
$\leq$ $\frac{\|\overline{w}^{(1)}_{c}(t)\overline{x}(t)\|^{2}{}\|\Delta\overline{x}(t)\|^{2}}{\|\tanh(\overline{w}^{(1)}_{c}(t)\overline{x}(t))\|^{2}}$
$=$ $\delta$
given the monotonicity of $y=\frac{|x|^{2}}{|\tanh(x)|^{2}}$. $\overline{\|\Delta x(t)\|}$ is the upper bound of $\|\Delta x(t)\|$ and determined by the system dynamics and the length of discrete time step.
To ensure negativity of the term 2:
$1-l_{c}\|\phi_{c}(t)\|^{2}-\frac{l_{c}}{\gamma_{2}}\|a(t)\|^{2}\|y(t)\|^{2}-\frac{1}{\gamma_{2}}>0$ (76)
Therefore,
$l_{c}<\min_{t}\frac{\gamma_{2}-1}{\gamma_{2}\big(\|\phi_{c}(t)\|^{2}+\frac{1}{\gamma_{2}}\|a(t)\|^{2}\|y(t)\|^{2}\big)}$ (77)
In particular, $\gamma_{2}>1$.
Similarly, to ensure negativity of the term 3:
$\frac{1}{\gamma_{1}}-\frac{2l_{a}}{\gamma_{1}}\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}-\frac{2l_{a}}{\gamma_{3}}\|C^{T}(t)D(t)\|^{2}\|x(t)\|^{2}-\frac{1}{\gamma_{3}}>0$ (78)
Therefore,
$l_{a}<\min_{t}\frac{1}{2}\frac{\gamma_{3}-\gamma_{1}}{\gamma_{3}\|C(t)\|^{2}\|\phi_{a}(t)\|^{2}+\gamma_{1}\|C^{T}(t)D(t)\|^{2}\|x(t)\|^{2}}$ (79)
In particular, $\gamma_{3}>\gamma_{1}$.
To ensure negativity of term 4:
$\frac{\lambda}{\gamma_{1}}-\frac{2\lambda^{2}l_{a}}{\gamma_{1}}\|U(t)\|^{2}\|\phi_{a}(t)\|^{2}-\frac{2\lambda^{2}l_{a}}{\gamma_{3}}\|U^{T}(t)D(t)\|^{2}\|x(t)\|^{2}-\frac{\lambda}{\gamma_{3}}>0$ (80)
Therefore,
$l_{a}<\min_{t}\frac{1}{2\lambda}\frac{\gamma_{3}-\gamma_{1}}{\gamma_{3}\|U(t)\|^{2}\|\phi_{a}(t)\|^{2}+\gamma_{1}\|U^{T}(t)D(t)\|^{2}\|x(t)\|^{2}}$ (81)
The fifth term can be bounded as
$\big\|w_{c}^{*(2)}\phi_{c}(t)-r(t)-\alpha\hat{w}^{(2)}_{c}(t)\phi_{c}(t+1)\big\|^{2}$ $\leq 4\|w_{c}^{*(2)}\phi_{c}(t)\|^{2}+4r^{2}(t)+2\alpha^{2}\|\hat{w}^{(2)}_{c}(t)\phi_{c}(t+1)\|^{2}$ (82)
Let $\overline{C},\overline{w}_{a1},\overline{w}_{a2},\overline{w}_{c1},\overline{\phi}_{a},\overline{y},\overline{x},\overline{a},\overline{D}$ be upper bounds of $C,w_{a1},w_{a2},w_{c1},\phi_{a},y,x,a,D$, and let $w_{c2}=\max\{w_{c}^{*(2)},\,w^{(M)}_{c2}\}$, where $w^{(M)}_{c2}$ is the upper bound of $\hat{w}^{(2)}_{c}(t)$. Finally, we obtain the following bound:
$\big\|w_{c}^{*(2)}\phi_{c}(t)-r(t)-\alpha\hat{w}^{(2)}_{c}(t)\phi_{c}(t+1)\big\|^{2}+4\|r(t)\|^{2}+\frac{8\alpha^{2}}{\gamma_{1}}\|w_{c}^{*(2)}(t)\phi_{c}(t+1)\|^{2}+\frac{1}{\gamma_{1}}\|C(t)\|^{2}\|\zeta_{a}(t)\|^{2}$ (83)
$\quad+\frac{1}{\gamma_{2}}\|a(t)\|^{2}\|\tilde{w}^{(1)}_{c}(t)y(t)\|^{2}+\frac{1}{\gamma_{3}}\|C^{T}(t)D(t)\|^{2}\|\tilde{w}^{(1)}_{a}(t)x(t)\|^{2}$
$\leq\Big(\frac{8\alpha^{2}}{\gamma_{1}}+4+2\alpha^{2}\Big)(\overline{w}_{c2}\overline{\phi}_{c})^{2}+8\overline{r}^{2}+\frac{1}{\gamma_{1}}(\overline{w}_{c(2)}\overline{C}\overline{w}_{a2}\overline{\phi}_{a})^{2}+\frac{1}{\gamma_{2}}(\overline{w}_{c1}\overline{ya})^{2}+\frac{1}{\gamma_{3}}(\overline{w}^{(2)}_{c}\overline{C}\overline{D}\overline{w}_{a1}\overline{x})^{2}=M$
The conclusion is, if $\gamma_{1}>8\alpha^{2}(1+\frac{\|L\Delta x(t)\|^{2}}{\|\phi(x(t))\|^{2}))}{\|\zeta(t)\|^{2}}$ in the term 1, and $l_{c}$, $l_{a}$ satisfying (77), 81, and $\|\tilde{w}^{(2)}_{c}(t)\|^{2}>\frac{M}{\|\phi(x(t))\|^{2}-\frac{8\alpha^{2}}{\gamma_{1}}\|\phi(x(t+1))\|^{2}}$, we obtain $\Delta L(t)<0$. Based on Theorem 1, this implies that the system of estimation errors is ultimately uniformly bounded.
V Online learning flight Control Design
In this section, angle-of-attack tracking control is designed using the TS-IHDP. To address the nonlinear aerodynamics that disturb control performance, a cascaded control structure is developed via a virtual control approach that explicitly incorporates the pitch-rate reference and is trained by two RL agents, namely the outer-loop and inner-loop agents. This architecture is also referred to as hierarchical learning. The switching pitch-rate reference is further passed through a low-pass filter to obtain smooth signals, as is commonly applied in command-filtered backstepping [Ref].
V-A Angle-of-attack tracking control
The dynamical model of aerial vehicles is given as [Ref]
$\dot{\alpha}$ $=\left(\frac{fgQS}{WV}\right)\cos(\alpha)[\phi_{z}(\alpha)+b_{z}\delta]+q$ (84)
$\dot{q}$ $=\left(\frac{fQSd}{I_{yy}}\right)[\phi_{m}(\alpha)+b_{m}\delta]$
where $\alpha$ is the angle of attack, $q$ is the pitch rate, $\delta$ is the control surface deflection. The aerodynamic coefficients are seen in Appendix. The tracking errors are defined as $e_{1}=\alpha-\alpha_{\text{ref}}$, $e_{2}=q-q_{\text{ref}}$, where $\alpha_{\text{ref}}$, $q_{\text{ref}}$ are reference signals, and
$\dot{e}_{1}=\dot{\alpha}-\dot{\alpha}_{\text{ref}}=f_{1}(\alpha)+e_{2}+q_{\text{ref}}+d_{1}-\dot{\alpha}_{\text{ref}}$ (85)
$\dot{e}_{2}=\dot{q}-\dot{q}_{\text{ref}}=f_{2}(\alpha)+g_{2}\delta-\dot{q}_{\text{ref}}$
The nonlinear functions are defined as
$f_{1}(x_{1})$ $=\left(\frac{fgQS}{WV}\right)\cos(\alpha)\phi_{z}(\alpha),\;d_{1}(x_{1},u)=\left(\frac{fgQS}{WV}\right)\cos(\alpha)b_{z}$ (86)
$f_{2}(x_{1})$ $=\left(\frac{fQSd}{I_{yy}}\right)\phi_{m}(\alpha),\qquad\quad g_{2}(x_{1},\delta)=\left(\frac{fQSd}{I_{yy}}\right)b_{m}$
The error dynamics 85 is a single-input-single-output (SISO) system with near-integrator dynamics. When pitch rate control is aggressive, the system exhibits oscillations. Moreover, the aerodynamics $f_{1}$ introduce disturbances to both the angle-of-attack tracking error and the control input.
V-B Incremental model
The incremental model of $q$-dynamics is
$\Delta q_{t+1}=F_{t-1}\Delta q_{t}+G_{t-1}\Delta\delta_{t}$ (87)
where the subscript $t$ denotes the discrete-time index. The state increments are defined as $\Delta q_{t+1}=q_{t+1}-q_{t}$, $\Delta q_{t}=q_{t}-q_{t-1}$, the control increment is defined as $\Delta\delta_{t}=\delta_{t}-\delta_{t-1}$. The first-order partial derivatives are $F_{t-1}=\frac{\partial f_{2}}{\partial\alpha}|_{\alpha_{t-1},\delta_{t-1},q_{t-1}},G_{t-1}=\frac{\partial g_{2}u}{\partial\delta_{t}}|_{\alpha_{t-1},\delta_{t-1},q_{t-1}}$. Only the signs of the control effectiveness $G_{t-1}$ are needed to construct policy gradients.
V-C Flight control design
V-C1 Cascaded control structure
The outer-loop control law and inner-loop control laws are parameterized by $q_{\text{ref}}=W_{1}(e_{1},\alpha,\delta)$, $\delta=W_{2}(e_{2},q,\alpha)$. The nested control law is
$\delta(e_{2},q,\alpha)=\underbrace{W_{2}(q-\underbrace{W_{1}(e_{1},\alpha,\delta)}_{\text{outer-loop actor}},q,\alpha)}_{\text{inner-loop actor}}$ (88)
where $W_{1}(\cdot;w^{A_{1}}),W_{2}(\cdot;w^{A_{2}})$ are neural networks with trainable parameter sets $w^{A_{1}},w^{A_{2}}$. The cascaded actor is illustrated in Figure 1, along with a comparison to the monolithic actor.
V-C2 Outer-loop agent
The outer-loop agent learns a pitch rate reference policy based on angle-of-attack tracking performance. Both the critic and actor are fully connected multilayer perceptrons (MLPs). The critic takes $\alpha,e_{\alpha},q]^{T}$ as input and outputs the estimated state value $\hat{V}_{1}$. The actor receives $\alpha,e_{1},\delta]^{T}$ as input, and its output is pitch rate reference, constrained using a scaled $\tanh(\cdot)$ function to enforce its limits. The critic is updated via temporal-difference (TD) learning to minimize $L^{C_{1}}_{t}=\frac{1}{2}\delta_{1(t)}^{2}$, where $\delta_{1(t)}=c_{1(t)}+\alpha\hat{V}^{\prime}_{1(t+1)}-\hat{V}_{1(t)}$ is the TD error, $\hat{V}^{\prime}_{1}(\cdot)$ is the target critic which delays weight updates to stabilize learning.
The one-step cost is
$c_{1}(t)=e_{1}^{2}(t)+aq_{\text{ref}}^{2}(t)$ (89)
The multi-objective optimization is formulated as
$q^{*}_{\text{ref}(t)}=\arg\min_{v_{t}}\Big(c_{1(t)}+\alpha\hat{V}^{\prime}_{1(t+1)}+\lambda_{1}L_{1(t)}^{\text{TS}}\Big)$ (90)
where the temporal smoothness loss
$L^{\text{TS}}_{1(t)}=\big|q_{\text{ref}(t+1)}-q_{\text{ref}(t)}\big|$ (91)
V-C3 Inner-loop agent
The inner-loop agent learns a control-surface deflection policy based the pitch-rate tracking performance. The non-minimum-phase dynamics associated with $\delta$ may inadvertently affect the $\alpha$-dynamics. This motivates the adoption of a cascaded control structure, which isolates this effect from the inner-loop agent’s training. As a result, the outer-loop agent is prevented from attempting to control the angle of attack directly using the control-surface deflection.
The critic receives $q,e_{q}]^{T}$ as input, and outputs the estimated state-value $\hat{V}_{2(t)}$. The actor receives $q,e_{q},\alpha]^{T}$ as input and outputs the control surface deflection, with scaled $\tanh(\cdot)$ at the output layer to enforce action limits. The critic is trained to minimize the loss $L^{C_{2}}_{t}=\frac{1}{2}\delta_{2(t)}^{2}$, where $\delta_{2(t)}=c_{2(t)}+\alpha\hat{V}^{\prime}_{2(t+1)}-\hat{V}_{2(t)}$ is the TD error and $\hat{V}^{\prime}_{2}(\cdot)$ is the target critic, updated as a delayed version of $\hat{V}_{2}(\cdot)$ to stabilize learning.
The one-step cost is
$c_{2}(t)=e_{2}^{2}(t)+bu^{2}(t)$ (92)
where $b\in\mathbb{R}_{+}$ trades off between tracking error and control effort. The optimal control law is defined as
The multi-objective optimization is formulated as
$\delta^{*}_{t}=\arg\min_{\delta_{t}}\Big(c_{2(t)}+\alpha\hat{V}^{\prime}_{2(t+1)}+\lambda_{2}L_{2(t)}^{\text{TS}}\Big)$ (93)
where the temporal smoothness loss
$L^{\text{TS}}_{2(t)}=\big|\delta_{t+1}-\delta_{t}\big|$ (94)
V-C4 Training
The action $q_{\text{ref}}$ is passed through the inner-loop actor to produce the control input $\delta$, which directly interacts with the plant model. The inner-loop actor therefore acts as an additional control effectiveness. From the perspective of the outer-loop actor, the plant is effectively an extended system that includes the inner-loop actor (see Figure 1). The outer-loop actor is trained using the transition sample $(e_{1(t)},\alpha_{t},q_{t},\delta_{t},q_{\text{ref}},e_{1(t+1)},\alpha_{t+1},q_{t+1})$, where the inner-loop actor is treated as part of the environment. In contrast, the inner-loop actor is trained independently of the outer-loop actor, since policy gradient propagation does not pass through the outer-loop policy. Its training sample is $(e_{2(t)},q_{t},\alpha_{t},\delta_{t},e_{2(t+1)},q_{t+1})$. Under this formulation, the environment for inner-loop training consists solely of the $e_{2}$-dynamics.
Remark 6.
Intuitively, the training of outer-loop actor may require the incremental model of $\alpha$-dynamics as it is based on the $\alpha$ tracking performance. The outer-loop actor still interacts with the model through control surface deflection, i.e. the $q$-incremental model, and the $\alpha$-incremental model is implicitly contained in $\hat{V}_{2}$.
Figure 1: Cascaded actor versus monolithic actor. The former uses an outer-loop actor to generate a pitch-rate reference, which is then passed to an inner-loop actor that produces the control-surface deflection, providing an explicit mechanism for learning and tracking the pitch-rate reference. The latter is a fully connected multilayer perceptron (MLP) that approximates an end-to-end tracking policy. The pitch-rate reference is not explicitly represented by the monolithic actor. This end-to-end structure is commonly used for the stabilization of inverted pendulum and Van der Pol’s oscillator, without additional angular dynamics.
V-C5 Low-pass filter
A second-order filter is used to smooth the pitch rate reference [Ref]:
$\dot{d}_{1}$ $=d_{2}$ (95)
$\dot{d}_{2}$ $=-2\zeta\omega^{2}_{n}d_{2}-\omega^{2}_{n}(d_{1}-q_{\text{ref}})$
where $d_{1},d_{2}$ are filtered signal and its differentiation, $\omega_{n}$ is the natural frequency, $\zeta$ is the damping factor.
Remark 7.
Policy training typically assumes a Markov environment, in which the one-step future state depends only on the current state. However, due to its integral dynamics, the filter state depends on the two previous states, thereby violating the Markov assumption when the filter is included during training of the outer-loop actor. Consequently, gradients from the value function $\hat{V}_{1}$ are delayed by two time steps. Although this issue can be addressed by augmenting the environment state with the filter states, doing so increases the computational burden because of the resulting higher system dimensionality. To eliminate gradient delay, this paper excludes the filter from backpropagation of the outer-loop policy gradient and applies it only during forward propagation.
Figure 2: Command-filtered cascaded online learning flight control system.
VI Simulation
Simulation results are presented for cascaded online-learning flight control systems. The aerial vehicle dynamics are numerically simulated through the fourth-order Runge Kutta method. The actuator dynamics are modeled as a first-order system $\tau\dot{\delta}_{t}+\delta_{t}=\delta_{c(t)},\tau=0.005$. To constrain actions, the outer-loop and inner-loop actors are scaled to $20^{\circ},20^{\circ}]$ and $20^{\circ}/\text{s},20^{\circ}/\text{s}]$. The reference of angle-of-attack is set as $\alpha_{\text{ref}(t)}=10^{\circ}\sin(\frac{2\pi}{T}t),T=10\text{s}$. The network weights are randomly initialized using a uniform distribution $\mathcal{U}(-0.01,0.01)$, which ensures an initially low-gain control law and slow learning to guard against inaccurate policy gradients. The persistence excitation signal is generated by adding Gaussian noise $\mathcal{N}(0,0.04)$ to the control surface deflection. The hyperparameters of the RL agents are listed in Table I. The baseline control system employs IHDP without a filter, i.e. $\lambda_{1,2}=0$. We discuss the performance in terms of tracking control, action smoothness, activation function saturation, action increments, and actor sensitivity.
TABLE I: Hyperparameters of RL agents
Parameter Outer-loop agent Inner-loop agent
critic learning rate $\eta^{C_{1}}$, $\eta^{C_{2}}$ 0.1 0.1
actor learning rate $\eta^{A_{1}}$, $\eta^{A_{2}}$ $5\times 10^{-7}$ $10^{-7}$
discount factor $\alpha$ 0.6 0.6
delay factor $\tau$ 1 1
policy iteration number at each time step 3 3
hidden layer size 7 7
critic hidden-layer activation function tanh tanh
critic output layer activation function abs abs
actor activation function tanh tanh
optimizer Adam Adam
control effort weights $a,b$ $5\times 10^{-6}$ $10^{-5}$
smoothness loss weight $\lambda_{1},\lambda_{2}$ $9.3\times 10^{-4}$ $10^{-5}$
loss thresholds for terminating critic training $2.5\times 10^{-4}$ $5\times 10^{-5}$
maximal steps of critic and actor updates 50 50
VI-A Tracking performance and control smoothness
VI-A1 Tracking control
In Figure 3, IHDP-based control system learns tracking the angle-of-attack reference during the first 10 seconds, then experiences growing oscillations with increasing tracking errors after 32s. The outer-loop and inner-loop actors are near-saturated after 32s and 23s, respectively. The amplitude of action oscillation is reduced by using a hard action constraint within $\pm 5^{\circ}$ in [Ref], but the policy becomes a bang-bang control with constrained amplitude. In contrast, TS-IHDP learns an action-increment-constrained policy, naturally generating actions within $\pm 6^{\circ}$, and prevents actor saturation as shown in Figure 7. In Figure 4, the critic weights converge to sub-optimal values within 10s. Oscillations happen after 30s in IHDP-based control system. The actor weights of TS-IHDP-based control system grow slightly more slowly due to the stabilized tracking errors.
Challenges on TS-IHDP-based control system. Control smoothness of TS-IHDP is sensitive to the empirical tuning of $\lambda_{1}$, $\lambda_{2}$. Large values reduce exploration of the state space, whereas small values degrade control smoothness. Moreover, TS-IHDP-based control system still exhibits frequently switching actions when tracking errors are near zero. This is because small action increments are only lightly penalized by the temporal smoothness loss (see Equations 91, 94). The frequently switching pitch rate reference further induces switching control surface deflections with amplified amplitudes through the inner-loop actor.
Therefore, we introduce a low-pass filter into the cascaded control system, which is commonly employed in command-filtered backstepping [Ref]. The natural frequency and damping ratio are $\omega_{n}=20\text{rad/s}$ and $\zeta=0.7$. The comparison between TS-IHDP and command-filtered TS-IHDP shows the filter generates a smoother pitch rate reference and further leads to smoother control surface deflection.
Figure 3: Tracking control performance.
Figure 4: Critic and actor parameters.
Figure 5: Spectrum of actions.
VI-A2 Control smoothness
Control smoothness is evaluated in the frequency domain using the Fast Fourier Transform (FFT), which transforms the action sequence from the time domain to the frequency domain, as defined by the Discrete Fourier Transform $X[k]=\sum_{n=0}^{N-1}x[n]\,e^{-j\frac{2\pi}{N}kn},k=0,1,\ldots,N-1$. The spectra in Figures 5 and 6 show that the components of $q_{\text{ref}}$ below 100 Hz are attenuated by TS-IHDP, and the signal amplitudes are significantly reduced without compromising tracking performance. Similar results are observed for control surface deflection. The switching control surface deflection is a result of the forward propagation of the switching tracking error $e_{2}$ through the inner-loop actor. For command-filtered TS-IHDP, these action signals in the 10-40Hz range are effectively attenuated by the filter.
Figure 6: Action smoothness in selected time intervals.
VI-A3 Activation function
The activation function $\tanh(\cdot)$ outputs values near its maximum or minimum, and its derivative vanishes when the inputs are large (see Figure 7). Consequently, policy learning becomes slow due to vanishing policy gradients. Another issue is the bang–bang control behavior of a saturated actor in response to small tracking errors, which degrades tracking performance and system stability.
Actor saturation results from an aggressive policy learning process, which occurs for two main reasons: (1) large learning rates. A large learning rate causes the actor weights to be updated with large increments, pushing the activation functions into their saturation regions. This process is effectively irreversible due to vanishing gradients. (2) Underemphasized control effort in the one-step cost function. A one-step cost function that underemphasizes control effort encourages the use of large actions during learning. As a result, the aggressive policy persists and continues to degrade tracking performance even as the tracking error decreases, due to vanishing policy gradients. During 10–20 s, switching actions occur even though $\tanh(\cdot)$ is not saturated, indicating that an unsaturated actor can also generate oscillatory actions. Moreover, the time-varying dynamics and time-varying signal $\alpha_{\text{ref}}$ cause switching tracking error $e_{\alpha}$, which propagate through the outer-loop actor and inner-loop actor to generated switching pitch-rate reference and switching control surface deflections. The pitch rate reference shows switching control without saturation, showing the actor does not need to be saturated to show bang-bang control behaviors
In Figure 7, IHDP makes the inputs of $\tanh(\cdot)$ within intervals [-4,-2], [2,4] (the outer-loop actor) and intervals [-3,-2], [2,3] (the lower-level actor). In these intervals, the derivative satisfied $\tanh^{\prime}(\cdot)\leq 0.1$ and attenuates the overall policy gradient. As a comparison, TS-IHDP and command-filtered TS-IHDP prevent reaching to the saturation regions. The inputs are kept within ’unsaturated’ intervals [-2,2] (the outer-loop actor) and [-0.5,0.5] (the lower-level actor). The derivative satisfies $\tanh^{\prime}(\cdot)\geq 0.4$ (the outer-loop actor) and $\tanh^{\prime}(\cdot)\geq 0.8$ (the inner-loop actor). This demonstrates the effectiveness of TS losses in mitigating actor saturation, as the saturated actor generates large action increments.
Figure 7: $\tanh(\cdot)$ in output layers of actors.
VI-A4 Action increment
The one-step action increments are $\Delta q_{\text{ref}(t+1)}=\left|q_{\text{ref}(t+1)}-q_{\text{ref}(t)}\right|$, $\Delta\delta_{t+1}=\left|\delta_{t+1}-\delta_{t}\right|$. Figure 8 shows the IHDP exhibits large-amplitude and fast-switching action increments, while TS-IHDP reduces the action increment to $|\Delta q_{\text{ref}(t+1)}|\leq 0.5\ ^{\circ}/\text{s}$ and reduces the amplitude of $|\Delta\delta_{t+1}|$.
Figure 8: Increments of actions.
VI-A5 Actor sensitivity
The actor sensitivity indicates how much, and in what direction, the output changes when the input increases slightly [Ref]. The sensitivity measures of the outer-loop and inner-loop actors are
$K_{1(t)}$ $=\frac{\partial q_{\text{ref}(t)}}{\partial e_{1(t)}}$ (96)
$K_{2(t)}$ $=\frac{\partial\delta(t)}{\partial e_{2(t)}}$
The relation between the pitch rate reference and its sensitivity measure is
$|q_{\text{ref}(t+1)}-q_{\text{ref}(t)}|$ (97)
$\approx$ $\left|K_{1(t)}(e_{1(t+1)}-e_{1(t)})\right|$
$=$ $\left|K_{1(t)}\right|\cdot\left|e_{1(t+1)}-e_{1(t)}\right|$
Similarly,
$|\delta_{t+1}-\delta_{t}|$ (98)
$\approx$ $\left|K_{2(t)}(e_{2(t+1)}-e_{2(t)})\right|$
$=$ $\left|K_{2(t)}\right|\cdot\left|e_{2(t+1)}-e_{2(t)}\right|$
Therefore, the actor sensitivity can be reduced through constraining action increments. In Figure 9, the comparisons on $K_{1},K_{2}$ show that applying a filter slows their growth and reduces high-frequency switches, contributing to a more robust controller.
Figure 9: Sensitivity measures.
VI-B Terminate and restart learning
VI-B1 Termination
The previous analysis assumes continual learning; that is, the agents never terminate learning during system operation. The issue lies in that growing control gains can cause state oscillations, even when tracking errors are small. Therefore, it is intuitive to terminate learning when the required tracking performance is satisfied. Specifically, the actor weights are fixed if the accumulated tracking error over a period $T_{s}$ remains within a specified threshold:
$\overline{e}_{s}\leq\epsilon$ (99)
where $\overline{e}_{s}=\frac{1}{n_{s}}\sum^{n_{s}-1}_{i=0}|e_{t-i}|$ is the average of absolute tracking errors, $n_{s}=T_{s}/\Delta T$ is the number of samples, $\epsilon>0$ is an error threshold. A large $\epsilon$ makes learning terminated early when tracking performance is unsatisfactory. Define the online-learning state indicator as $\tau\in{1,0}$, with $\tau=1$ denoting that learning is off and $\tau=0$ denoting that learning is on.
During the initial phase, when $t<T_{\text{s}}$ and there are not enough samples to compute $\overline{e}_{\text{s}}$, learning remains active:
$\tau=\begin{aligned} &0,\;\text{if}\;\;t<T_{\text{s}}\end{aligned}$ (100)
Figure 10: Termination case (left) and restart case (right). The shaded area represents a performance threshold. The actual average tracking errors must remain below this threshold to terminate learning, and exceed it to trigger a restart.
VI-B2 Restart
The learning must restart when the historical control performance degrades, typically occurring when the control system falls into faults and uncertainties. The restart condition determines in what case to restart learning. The evaluation period, denoted by $T_{r}$, should be carefully chosen to ensure it qualifies to represent the required control performance. A long evaluation period would delays the restart, preventing a timely response to performance degradation, while a short evaluation period makes the restart decision overly sensitive, potentially triggered by measurement noise or short-term peaks of tracking errors. As a result, the learning state may switch frequently between on and off.
The restart condition is
$\overline{e}_{r}\geq\sigma$ (101)
where $\overline{e}_{r}=\frac{1}{n_{r}}\sum^{n_{r}-1}_{i=0}|e_{t-i}|$ is the averaged absolute tracking errors, $n_{r}=T_{r}/\Delta T$ is the number of samples, $\sigma>0$ is an error threshold. The sketch maps are provided in Figure 10. The pseudocode is seen in Algorithm 1.
The outer-loop and inner-loop agents have independent termination and restart conditions. Their parameters are seen in Table II. Figure 11 shows the histories of average tracking errors and the indicators. The outer-loop agent is turned off at 29s and restarted at 31.5s, while the inner-loop agent is turned off at 22s and is never restarted.
Data: $T_{s},\epsilon,T_{r},\sigma$, $t$, $\Delta T$
1 Initialization termination indicator $\tau_{-1}=0$
2 for each environment step $t\leftarrow$ to $t_{\text{max}}$ do
3 if $\tau_{t-1}=0$ then
4    if $t\leq T_{s}$ then
5       $\tau_{t}=0$
6    else
7       if $\overline{e}_{s}\leq\epsilon$ then
8          $\tau_{t}=1$
9       else
10          $\tau_{t}=0$
11         end if
12
13      end if
14
15   end if
16
17 end for
18if $\tau_{t-1}=1$ then
19 if $\overline{e}_{r}\geq\sigma$ then
20    $\tau_{t}=0$
21 else
22    $\tau_{t}=1$
23   end if
24
25 end if
Algorithm 1 Termination and restart algorithm
TABLE II: Parameters
Parameter Outer-loop agent Inner-loop agent
$T_{s}$ 10s 10s
$\epsilon$ 0.25 0.5deg/s
$T_{r}$ 0.2s 0.1s
$\sigma$ 0.8 2deg/s
Figure 11: Termination and restart measures.
VII Conclusion
We studied temporally smoothed incremental heuristic dynamic programming (TS-IHDP) for online learning. Trace analysis shows that the critic and actor weights converge to suboptimal values in the sense of ultimate uniform boundedness (UUB), where the actor is updated using an additional temporal smoothness loss. Flight control simulations verify the effectiveness of TS-IHDP in learning an action-increment-constrained policy, thereby avoiding actor saturation and degraded tracking performance. The low-pass filter is shown to effectively remove high-frequency components of the pitch-rate reference, which cannot be achieved by TS-IHDP alone. The combination of these two methods enables a sufficient level of action smoothness during online learning. These techniques are essential for online learning-based tracking control of aerial vehicle systems with time-varying dynamics, which are prone to state and action oscillations caused by switching tracking errors.
Appendix
VII-A Aerodynamic coefficients
The aerodynamic coefficients are approximately computed by $b_{z}=-0.034,b_{m}=-0.206$, and
[…]
These approximations of $b_{z},b_{m},\phi_{z}(\alpha),\phi_{m}(\alpha)$ hold for $\alpha$ in the range of $\pm 20$ degrees. The physical coefficients are provided in Table III. In addition, the actuator dynamics are considered as a first order model with the time constant $0.005\text{s}$. The rate limit is 600 deg/s, and a control surface deflection limit is $\pm 20$ degrees.
TABLE III: Physical parameters (adapted from [bib32])
Notations Definition Value
$g$ acceleration of gravity 9.815 m/s2
$W$ weight 204.3 kg
$V$ speed 947.715 m/s
[…]
$Q$ dynamic pressure 29969.861 kg/m2
$S$ reference area 0.041 m2
$d$ reference diameter 0.229 m